A Complete Bibliography of Publications in the ACM Journal of Experimental Algorithmics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 December 2023 Version 1.27

Title word cross-reference

(n log n + 0.2n) [ES02]. (n log n − 0.9n) [ES02]. 1 [ABC^{+}18, LNW08]. 2 [DDF19, DFKM09, LMMM05]. b [MHS99, MHS00]. k [BHM^{+}19, CKLM09, CAGDPS^{+}23, EP16a, EP16b, FR15, GKN17, HT17, NW13, SF18, ACN23]. n [AKK19]. n × m [Shi00]. O(nm log n) [MS02]. q [STK06]. R [HW11]. R^{3} [DFKM09]. t [FG09]. Z [SF18].

-cardinality [CKLM09]. -center [CAGDPS^{+}23, LNW08]. -fold [AKK19]. -grams [STK06]. -Hop [DDF19]. -manifold [DFKM09]. -matching [MHS99, MHS00]. -Median [NW13]. -Nearest [EP16a, EP16b]. -Order [SF18]. -Plexes [BHM^{+}19]. -spanners [FG09]. -trees [HW11].

2005 [DT08]. 2006 [SÁ09]. 2008 [MW09]. 2010 [Fes13]. 2013 [SZ16a, SZ16b]. 2014 [GK16]. 2017 [PC18]. 2018 [D’A19]. 2022 [FKKS23].

32- [NG10].

64-bit [NG10].

Accelerating [VBC18]. accessing [YZ98]. aCGH [TPT^{+}11]. acyclic [PK06]. Adapting [RR01]. Adaptive [IGA05, MPR20, PF08, BN19]. adaptiveness [BFM08]. Adding [BCD^{+}18]. Address [JM15]. Advanced [SSS15]. Advances [HHS22]. against [BSTU08]. air [BMS09]. airspace [BMS09]. ALENEX [DT08, MW09, PC18, SZ16a, SZ16b]. ALENEX’12 [BM15]. Algorithm [ABC^{+}18, ABM22, BLW16, BN19, CEF23, CMN16, DPSW18, GC20, JB20, KBJ19, NG22, RMH^{+}16, BDS^{+}10, BLMM05, BGL03, BFV08, BS99, CEM13, CT09, CW09, EJ99, FT06, GS08, HJB98, Kim99, MPR04, MSS^{+}06, PK06, SWW00, VSM03, YZ98]. Algorithmic [BLN22, BG97]. Algorithms [AM20, ATV01, BM19, BDSU22, BKP20, BM21, CFP19, CHW19, DDF^{+}15, EKX09, FFF15, FMZ01, GD15, GL06, HHS22, HNSS18, IKRT01, KKP16a, KKP16b, LW22, MO15, MNG16, MW23, MNNW15, NPR^{+}19, NW13, PS06, PW17, Sto19, dJvdLT23, AS02, ACI97, BHJ96, BLOLS09, BMS09, BKL00, CGM^{+}98, CGG^{+}09, DL10, DK05, EJ02, ECHS09, FINP98, GGHN08, HRSZ98, HSW^{+}09, JMN99, Jac10, KM13b, KZ08, Li08, Mag98, MM06, Mic11, MHS99, MHS00, PCJ97, RN11, SRB08, TPT^{+}11, Ull10, XZK00]. Alignment [DLY23, QSL^{+}21]. all-against-all [BSTU08]. all-pairs [VSM03]. allocation [BBJP08]. Alternative [ADGW13, LS15]. Altruistic [MO15]. amalgamated [GT04]. Anagrams [Rea12]. Analysing [RR00]. Analysis [GK01, LLH^{+}22, RVHE18, Boy11, DFKS11, FINP98]. Analytical [DL10]. analyzed [BLMM05]. Anatree [Rea12]. Annealing [CEF23, JLM^{+}21]. Anonymity [dJvdLT23]. apart [CNV22]. Applications [CILP20, Epp00]. Applied [CDF^{+}23]. apportionment [HRSZ98]. Approach [DLY23, BSTU08, DS13, MP08]. approaches [AN10, CMP^{+}08, LMMM05]. Approximate [HT17, KMY03, Boy11, CEM13, FMRT02, FN04, HFN05, SZ05]. Approximating [ABM22, DSW19, FR15, PPR05, SMEDM08]. Approximation [NPR^{+}19, PW17, TPT^{+}11, BLMM05, DK05, EJ02, FCNP09, FNP06, LNW08, NR08, BN19]. Approximations [BC19]. Arc [BSNA21, GJR22]. Area [PPM16, IJS^{+}06]. arrangement [Pet03]. arrangements [HH08]. Array [CILP20, KK16, KK19, KM17, DKMS08]. Arrays [BFO16a, BFO16b, GNF15]. aspects [BG97]. Assignment [BSW19, KST20, EJ99, Jac10, Li08]. Audit [GD15]. Augment [CGM^{+}98]. automata [NR00]. Automated [LF19]. Automorphism [Sto19]. Autotopism [SFKM20]. Average [FN04, Li08]. average-case [Li08]. Average-optimal [FN04]. Avoiding [EW19].

B [ADT03]. B-trees [ADT03]. backtracking [GS08]. backtracking-based [GS08]. balance [HRSZ98]. balancing [BMS09]. BALL [BKL00]. balls [KMY03]. barriers [BA06]. Based [ATV01, CFP19, GK01, GMS16, HNS20, HSS19b, KM17, MRS01, NG22, AN10, Bel23, CKM00, CEF23, CS00, DHW08, FT06, GS08, NAH04, QSL^{+}21]. basis [BW02, MM06, PCJ97]. Batch [ABM22, DSW19]. Batch-dynamic [ABM22]. Batches [DDF^{+}15]. Benchmark [FGLP21, HMP^{+}18]. Benchmarking [BGQ19]. Benchmarks [Spe15, Spe10]. best [DT09]. Better [DKMS08, KK19, KST20]. between [BM21, CMTW19, SMEDM08]. Betweenness [BCD^{+}18, BN19, VBC18]. biased [CS98]. bidirected [EJ02]. Bigraph [SBG01]. Billiard [CEF23]. bin [BBJP08]. Binary [CDF^{+}23, GL22, DS13, Ull10]. biobjective [DS13]. biochemical [BKL00]. Bipartite [BKP20, CGM^{+}98, LMS10, NPS^{+}11]. Bipartization [GHS21]. Bisection [HS18]. Bit [HT17, Ull10, HFN05, NG10, NR00]. Bit-Parallel [HT17]. bit-parallelism [HFN05, NR00]. Bit-vector [Ull10]. blocked [VSM03]. BlockQuicksort [EW19]. Bloom [GL20, PSS09]. Boltzmann [AON15]. Boolean [ARMS08]. boosted [SS23]. Bose [AON15]. Bouchitté [KBJ19]. Bound [CMN16, FT06]. Bounded [AGG^{+}23, ZP19, BFK^{+}03, Jul09]. Bounded-Degree [AGG^{+}23]. bounds [dSLAM05]. Branch [BLW16, CMN16, EW19, BNWG08, FT06]. Branch-and-Bound [CMN16, FT06]. Branch-Price-and-Cut [BLW16]. branching [BLLW13]. Breaking [DF01, ARMS08]. Broadcast [Lib01]. broadcasting [FNP06]. BST [AZ10]. Buffered [FS22]. buffers [ERW09]. build [SZ05]. burst [AZ10]. burstsort [SW10]. bypassing [YZ98].

Cache [BM21, FPR09, PP06, PSS09, SZ04, SZR06, AZ10, BFV08, ERS99, RR00]. Cache- [PSS09]. Cache-conscious [FPR09, SZ04]. Cache-efficient [SZR06]. Cache-Friendly [PP06]. cache-oblivious [BFV08]. cached [San00]. caches [LL96, WACV02]. Caching [Lib01]. Candidate [LS15]. Canonization [AM20]. Capacitated [CAGDPS^{+}23]. capacity [CGM^{+}98]. Cardinality [DW15, Spe15, CKLM09]. case [BKS00, ERS99, LRAM06, Li08, SWW00]. center [CAGDPS^{+}23, LNW08]. Centrality [BCD^{+}18, VBC18]. CG [FKKS23]. CGAL [FHH^{+}00, HH08]. Challenge [FKKS23]. checkpointing [BA06]. Chips [GC20]. chordal [PPR05]. chromatic [HH02]. Circular [HT17]. class [SOS05]. class/teacher [SOS05]. Clique [FFF15, GGHN08]. Cliques [ELS13, KBJ19]. closest [Epp00]. Closure [FMZ01, KZ08, PP06]. cluster [MSM09]. ClusterHulls [HSS08]. Clustering [GKW15, Ano08, Epp00, GMS^{+}13, RN11]. Clusterings [CMTW19]. Clusters [CMTW19]. CNF [EOP20, IKM^{+}02]. Coarsening [GMS16, SSS15]. Collapse [DSW19]. Collections [NP16]. Collisions [NG22]. Colored [LPR21]. Coloring [BM19, CDF^{+}23, Man18, NPR^{+}19, SS23, CEM13, MPR04, PPR05]. Column [AB15, CGM09]. Combining [BDS^{+}10, HSWW05, GMS^{+}13, NR00]. Comments [GJR22]. common [BGL03]. communication [BHJ96].

Compact [BBG^{+}17, GKN17, NPV15]. Comparative [Boy11, CZ15]. Compare [KMN17]. Comparing [CMTW19, NG10]. Comparison [FPR23, JLM^{+}21, KM17, DL10, Mic11]. Comparison-Based [KM17]. comparisons [ES02]. Competition [FHH21]. Complex [BGQ19, GMS16, dJvdLT23]. Complexity [Man18, dSLAM05]. Compressed [GO13, GNF15, GO15, NP16, AN10, VMGD09]. Compression [MRS01, NLG^{+}22, LOMSS05]. Computation [BBDW19, CNV22, DHLO09, KLC15, vdBG^{+}22, GO13, WWZ05]. computational [CGM^{+}98, JMN99, MHS00]. computationally [HLL06]. computations [BSWW04, HSWW05]. Computing [BM19, BM21, CCL15, CMN16, JLM^{+}21, LNW08, PGI18, SSMJ99, Shi00, SFKM20, TSP18, WEM11, dJvdLT23, HTVW08]. concept [Ner02]. concurrently [YZ98]. Conflict [CDF^{+}23]. congestion [PS06, WSdC00]. Congress [HRSZ98]. Congressional [LF19]. Connection [DPSW18]. Connectivity [IKRT01]. conscious [FPR09, SZ04]. consideration [NTB05]. constants [TMH11]. Constrained [DPW15, GKW15, CEM13, TRC11]. Constrainedness [EOP20]. Constraint [DW15, Ull10]. Constraints [Spe15]. Constructing [SF18]. Construction [BBO17, DEF^{+}21, KK16, KK19, BKS00, DKMS08]. Constructive [CAGDPS^{+}23]. contact [IJS^{+}06]. containers [WWZ05]. Contraction [BSW19, DSW16, BGSV13]. controller [BMS09]. convex [BLLW13]. cooperative [Ner02]. coordinated [BA06]. coordinates [BSWW04]. copying [SZR06]. core [KCC11, KMY03]. core-sets [KMY03]. Correspondence [CS00]. Correspondence-based [CS00]. cost [HTVW08]. costs [ADGM06]. Counting [FHH21, FFF15, JB20]. couples [BIS11]. Cover [DDF19, DL10, FT06, GGHN08]. covering [ECHS09, FT06]. covers [HLL06]. Creates [Spe15]. Critical [PGI18, HH02]. Crossing [RRRW19, SBG01, CGM09, CGMW10]. crossings [DF01]. CSP [CDF^{+}23]. Cuckoo [GL20, MPR20]. Curve [SF18]. curves [HW11]. Customizable [BSW19, DSW16]. Cut [BLW16, HNSS18, MW23]. Cuts [BLW16]. cutting [Lev00]. Cycle [FEMPS16a, FEMPS16b, ZP19, MM06]. cycles [DF01].

D [ABC^{+}18, LMMM05]. Data [BDSU22, DLY23, FPGI13, GGHN08, KKN^{+}21, MRS01, Rea12, CS00, ERS99, HSS08, IGA05, MS02, NG10, NAH04, TPT^{+}11, VH02]. Data-dependent [DLY23]. datapath [dSLAM05]. David [Mcg16]. Dealing [TMH11]. decomposed [KKT^{+}20]. decomposition [GS08]. Decompositions [GLLO22, GO15, NZ01]. Decreasing [CNV22]. Degree [AGG^{+}23, Ull15]. Delaunay [KCC11]. delay [CEM13]. deliver [ECHS09]. denoising [TPT^{+}11]. Density [GKW15]. Density-Constrained [GKW15]. Dependent [KLC15, BGSV13, DLY23]. Design [CKM00, GK01, LLH^{+}22, BMS09, FHH^{+}00, MHS99]. Detection [LLH^{+}22]. Deterministic [BM19, HJB98]. Diagrams [BBG^{+}17, KMN17]. diameter [Jul09]. dictionaries [FPGI13]. Dictionary [MRS01, Boy11, DHW08].

Dictionary-Based [MRS01]. different [CMP^{+}08]. Difficult [EOP20, Spe10]. Digital [JLM^{+}21]. Dijkstra [BDS^{+}10, MSS^{+}06, SWW00]. Dimensional [SF18, DLY23, HW11, KM13b, SSMJ99]. dimensions [KMY03]. Directed [CMN16, PGI18, BDS^{+}10, EJ99, MSM09, PK06]. Direction [LLH^{+}22]. Direction-optimizing [LLH^{+}22]. Discovery [LPR21]. discrepancy [PU11]. disjoint [EJ02]. Disk [Lib01]. Distance [BM21, CNV22, SMEDM08]. distances [MSM09]. Distributed [BM19, MPR04]. Distribution [NG22, GL06, RR00]. Distributions [AON15]. diversification [SOS05]. Document [GKN17, NPV15]. Donation [MO15]. double [CS00, DT09]. double-ended [CS00]. double-tree [DT09]. drawing [PCJ97]. Droplet [GC20]. DSM [BA06]. dual [BLMM05, dSLAM05]. dummy [NTB05]. Dynamic [BM19, BLN22, DDF^{+}15, DDF19, FCLMC21, FMZ01, GMS^{+}13, HHS22, IKRT01, KKT^{+}20, MRS01, NR08, TW09, VBC18, VM18, vdBG^{+}22, ACI97, ABM22, DS13, DHW08, Epp00, FINP98, GL06, Hof13, KZ08, LFLSW08, PK06, Rad98, SZ04, TPT^{+}11].

’Easy [FMRT02]. Edge [BM19, EJ02, GRST12, MPR04, TRC11]. edge-coloring [MPR04]. edge-disjoint [EJ02]. editor [FH11]. Editorial [GK16, Kla15, PC18, Zar19]. Effect [MRS01]. Effective [KCC11]. effects [RR00]. Efficient [ABC^{+}18, ATV01, AB15, BBO17, CFP19, DSW19, EJ99, GO13, GC20, HMP^{+}18, KLC15, MNNW15, PF08, PSWZ08, TS16, WACV02, ADT03, ECHS09, HH11, MP08, NTB05, PSS09, SOS05, SZ05, SZR06, WWZ05, YZ98]. Efficiently [dJvdLT23, Kim99, Shi00]. Einstein [AON15]. elevation [WEM11]. Elimination [QSL^{+}21]. ELRUNA [QSL^{+}21]. Embeddings [BFKK20]. Empirical [FGLP21, NPR^{+}19, PPM16, ACI97, BIS11, SWW00]. enclosing [KMY03]. ended [CS00]. energy [FNP06]. Engineered [BSW19]. Engineering [AM20, Ano08, ACF^{+}21, BFV08, HSW08, HSW^{+}09, MNNW15, SW10, VMGD09, BKS00, TMH11]. enhanced [ERS99]. enough [DT09]. enumerating [CEM13]. Enumeration [CNV22, RMH^{+}16]. equilibria [PS06]. equilibrium [CMP^{+}08]. ESA [Zar19]. estimating [IJS^{+}06]. Estimation [CEF23]. Evaluating [AKK19]. Evaluation [CMN16, ERW09, GNR16, GHS21, KBJ19, MW23, NPR^{+}19, NW13, PW17, SBG01, VBC18, ZP19, Ano08, CGG^{+}09, HJB98, LRAM06, Li08, dSLAM05]. evolution [Ner02]. evolutionary [SMEDM08]. Exact [BSNA21, BBDW19, CHW19, GJR22, KM13b, Sto19, CGM09, GGHN08]. Expected [TSP18]. Experimental [BKP20, CMN16, FG09, FPR23, FR15, FINP98, FMZ01, GK01, GHS21, IKRT01, JLM^{+}21, KBJ19, Li08, LLH^{+}22, Man18, MW23, NW13, PW17, SBG01, ZP19, AS02, Ano08, BLOLS09, BNWG08, BCFM00, CGG^{+}09, CMP^{+}08, DL10, DFKS11, FPGI13, HH08, HJB98, Jac10, KZ08, LG02, Mag98, MPR04, Mic11, PCJ97, SSMJ99]. Experimentation [MO15]. Experiments [CGM09, CHW19, DDF^{+}15, FFF15, FMM20, GKW15, KKN^{+}21, Pet03, VV00]. explicit [VV00]. exploit [AZ10]. Explorable [FMM20]. exponential [SHA97]. Extending [EP16a, EP16b]. Extensions [JB20]. External [BFO16a, BFO16b, KK16, KK19, BCFM00, DKMS08]. Extremal [MNG16].

Factor [CFP19, FNP06]. Factorization [KKP16a, KKP16b]. Far [CNV22]. Far-apart [CNV22]. Fast [AW09, BD09, BBDW19, BFK^{+}03, DT09, Epp00, GL22, GO15, LOMSS05, NR00, PW17, Rea12, San00, BBP11, CKLM09, SW10]. Faster [BGQ19, CNV22, GL20, KM13a, LFLSW08, NP16, SSS21, FT06]. faults [FPGI13]. feasibility [CGG^{+}09]. Feedback [BSNA21, GJR22, HGH21]. Filters [GL20, GL22, MPR20, PSS09]. filtration [DSW19]. Finding [HH02, IM09, Lev00, MNG16, WSdC00, ZP19, ACN23]. Findings [FGLP21]. Fingerprinting [Bel23]. Fingerprinting-based [Bel23]. fingers [DM05]. Flexible [MRS01, GRST12, MHS99, NR00]. Flow [BBG^{+}17, GK01, HSS19b, CGM^{+}98, DK05, NPS^{+}11]. Flow-Based [HSS19b]. fold [AKK19]. Following [HMP^{+}18]. Foreword [Fes13, FH11]. Formula [BGQ19]. formulas [ARMS08]. Formulations [CZ15]. Four [HW11]. Four-dimensional [HW11]. FPT [ECHS09]. FPT-algorithms [ECHS09]. Framework [AM20, LLH^{+}22, LRAM06]. free [CGMW10]. Friendly [PP06]. Frontiers [VBC18]. Fully [BM19, BLN22, DDF19, HHS22, IKRT01, KZ08]. function [SHA97]. Fuse [GL22].

games [PS06]. Gauss [WEM11]. General [NPV15, MS02]. Generalized [GLLO22, NPR^{+}19]. Generating [BSWW04, EOP20, TRC11]. Generation [HMP^{+}18, CGM09]. generator [Spe10]. Generic [AM20]. Genome [GD15]. genomes [SMEDM08]. Geometric [AGG^{+}23, BMS09, NZ01, RRRW19, TSP18, WWZ05, FG09]. Geometry [KMN17, DFKM09]. Goal [MSM09, BDS^{+}10]. Goal-directed [MSM09, BDS^{+}10]. good [AW09, DT09]. GPU [CT09, RVHE18]. GPU-Quicksort [CT09]. Grammar [NLG^{+}22]. grams [STK06]. Graph [AM20, BM19, CDF^{+}23, FS22, GHS21, GKW15, HS18, HHS22, HNS20, KBJ19, NPR^{+}19, SSS15, SF18, Sto19, Ull15, ACI97, Ano08, BSTU08, GMS^{+}13, NTB05, PCJ97]. graphics [CT09]. Graphs [AB15, BHM^{+}19, BLW16, ELS13, FEMPS16a, FEMPS16b, HMP^{+}18, KKN^{+}21, LLH^{+}22, Man18, PGI18, SS23, VM18, ZP19, vdBG^{+}22, ADGM06, HH02, HSW08, HH11, MS02, MSS^{+}06, NPS^{+}11, PK06, PPR05, TRC11]. GRASP [FPPR06]. Greeding [Mag98]. Greedy [BFKK20, Jul09, BLMM05, EJ99]. Grid [ATV01, BS99, HTVW08]. Grid-Based [ATV01]. Gromov [CCL15]. Group [Sto19]. Groups [SFKM20]. Guest [FH11]. Guide [HHS22].

Hafnian [BGQ19]. Hamiltonian [ZP19]. ’Hard [FMRT02]. Harder [Spe15]. Hardest [Man18]. hash [AZ10, PSS09, SHA97]. hash- [PSS09].

Hashing [Bel23, BBPV11, DHW08, Kim99]. heap [BKS00]. heaps [LL96, YZ98]. HEAPSORT} [ES02]. Helps [KMN17]. Heuristic [CAGDPS^{+}23, LMS10, Sto19, BP97, FNP06, LMMM05, SOS05, dSLAM05]. Heuristics [BW02, IJS^{+}06, RRRW19, SBG01, BBJP08, CKM00, FMRT02, Jul09, NTB05]. hidden [TMH11]. Hierarchical [NW13, BDS^{+}10, Epp00]. Hierarchies [BSW19, DSW16, BGSV13]. Hierarchy [RR01]. High [DLY23, RVHE18, SHG^{+}22, SF18, KMY03]. High-Dimensional [SF18, DLY23]. High-Performance [RVHE18]. High-Quality [SHG^{+}22]. Hilbert [HW11]. History [GK01, BP97]. history-sensitive [BP97]. Homogeneous [DDF^{+}15]. Hop [DDF19]. hotlink [Jac10]. Hub [EP16a, EP16b]. Hybrid [HNR^{+}99, DS13]. hydrogens [LFLSW08]. HyperBench [FGLP21]. Hyperbolic [BFKK20, VM18]. Hyperbolicity [CCL15, CNV22]. Hyperedge [BDSU22]. Hypergraph [HSS19b, SHG^{+}22, CKM00]. Hypergraphs [FGLP21]. Hypertree [GLLO22, GS08].

I/O [ATV01, ADT03, HMP^{+}18]. I/O-Efficient [ATV01, HMP^{+}18, ADT03]. ILP [HNS20]. ILP-Based [HNS20]. Implementation [EJ02, LLH^{+}22, MS02, Rad98, CKM00, DHW08, EJ99, FHH^{+}00, HH11, VMGD09]. implementations [PP06]. Implementing [AN98, DK05, ES02, MM06, MHS99, MHS00, TS16]. Importance [JB20]. Improved [LMP23, KM13a]. Improving [BCD^{+}18, DM05, XZK00, NAH04]. in-place [SW10]. in-subtree [Rad98]. Increased [HFN05]. Incremental [GLLO22, NW13]. Independent [BLN22, HSS19a]. Index [ABC^{+}18]. Indexes [GKN17]. Indexing [BCF^{+}21, Boy11, CILP20]. indices [AN10]. Induced [NLG^{+}22]. Inducing [BFO16a, BFO16b]. influence [LL96]. information [PSWZ08]. initialization [LMS10]. insights [MHS00]. Inspired [BM19]. Instances [EOP20]. instruction [VV00]. Integer [AKK19, BHJ96, BLLW13, NG10]. Integrating [BA06]. interactive [LMMM05]. Intersection [AB15, BLOLS09]. interval [CEM13, PPR05]. intervals [Knu96]. Introduction [BM15, D’A19, SZ16a, SZ16b, BG97]. investigation [BLOLS09]. IP [IGA05, PF08]. Irredundant [Knu96]. Isocontours [BBDW19]. Isomorphism [Sto19, Ull10]. Issue [BM15, D’A19, GK16, PC18, SZ16a, SZ16b, Zar19, Fes13].

JEA [Nik06]. Johnson [Mcg16].

KADABRA [BN19]. Katz [vdBG^{+}22]. Kernelization [HSS19a]. key [GL06]. keys [NG10]. Kidney [MO15]. knapsack [DS13].

Label [LLH^{+}22]. Labeled [Ull15]. Labeling [BLN22, DDF19, GNR16, BLMM05]. Labels [EP16a, EP16b]. Large [EP16a, EP16b, ELS13, vdBG^{+}22, IM09, SZ04, TMH11, VH02]. Large-Scale [EP16a, EP16b]. Latin [SFKM20]. lattice [BW02]. Layer [CGMW10]. Layer-free [CGMW10]. layered [IJS^{+}06]. layering [NTB05]. layout [NAH04]. Layouts [KM17]. Lazy [KKP16a, KKP16b]. LCP [BFO16a, BFO16b, GO13, KK16, KK19]. LCP-values [GO13]. learning [BGL03, Ner02]. least [HTVW08]. least-cost-path [HTVW08]. leftist [CS98]. Lempel [AN10, ACF^{+}21, KKP16a, KKP16b]. level [AAS^{+}19]. LFR [HMP^{+}18]. library [BKL00]. lifetime [EKX09]. Limited [PU11]. line [SWW00]. Linear [CFP19, JB20, LMP23, NG22, Pet03, SRB08]. Lines [DLL23, ECHS09]. Links [BCD^{+}18]. list [LG02, VV00]. Listing [BHM^{+}19, ELS13]. Lists [FPR09, CS98]. Local [AW09, CW09, HNS20, IKM^{+}02, RN11]. locality [NAH04]. localization [LRAM06]. Localizations [HGH21]. Locally [GNF15]. location [ADT03, HH08, VH02]. Lock [YZ98]. Long [LW22]. Long-tailed [LW22]. longer [EKX09]. longest [BGL03]. lookup [PF08]. lookups [IGA05]. LP [AW09].

Machines [SSS21]. Maintenance [DDF^{+}15, BS99]. manifold [DFKM09]. manufacturing [IJS^{+}06]. Map [BLN22]. Mapping [KST20, KM13a]. MapReduce [FFF15]. Maps [GNR16, FHH^{+}00]. market [CMP^{+}08]. Massive [HMP^{+}18, SS23, HTVW08]. Matching [ABM22, AB15, BM19, BKP20, CFP19, HT17, KKN^{+}21, BSTU08, BIS11, BFK^{+}03, CGM^{+}98, FN04, HFN05, Kim99, KM13b, LMS10, Mag98, Mic11, MHS99, MHS00, NR00, STK06]. Matchings [CMTW19, FMRT02, IM09, MS02]. Matrices [BGQ19]. Matrix [ERS99]. matroid [DHLO09]. Max [MW23, PPR05, BP97]. max-coloring [PPR05]. MAX-SAT [BP97]. maxima [WEM11]. Maximal [BM19, BHM^{+}19, ELS13, KBJ19]. maximization [Li08]. Maximum [ABM22, CHW19, HSS19a, KKN^{+}21, PW17, FMRT02, NPS^{+}11]. MAXSAT [FPPR06]. Maxwell [AON15]. means [HH02]. Measures [TSP18]. mechanisms [BA06]. Median [NW13]. memetic [BS99]. Memoriam [Mcg16]. Memory [BFO16a, BFO16b, KK16, KK19, RR01, BBJP08, BCFM00, DKMS08, FPGI13, San00, XZK00]. Merging [VBC18, dSLAM05]. mesh [KCC11]. Method [BSNA21, GJR22]. Methods [GHS21, Boy11, HNR^{+}99]. metric [BLMM05, DT09, FCNP09]. Microfluidic [GC20]. Minimal [Bel23, BBPV11]. Minimization [RRRW19, SBG01, CGM09, CGMW10]. minimizing [DF01]. Minimum [BSNA21, BGSV13, CDF^{+}23, FKKS23, FMM20, GJR22, HNSS18, NZ01, DT09, FNP06, Jul09, KMY03, MM06, NTB05, Pet03, WSdC00]. minimum-weight [DT09]. minimum-width [NTB05]. Mismatches [HT17]. Mispredictions [EW19]. mixed [BLLW13]. mixed-integer [BLLW13]. Modal [KLC15]. Model [FHH21, JM15]. modeling [BKL00]. Models [CHW19, PSWZ08, Ano08]. modified [CS98]. modularity [GMS^{+}13, RN11]. molecular [BKL00]. monotone [BBPV11]. motivate [HH11]. Motorcycle [HH11]. move [CKM00]. move-based [CKM00]. MST [FNP06]. much [DFKM09]. Multi [AAS^{+}19, KLC15, VV00]. Multi-level [AAS^{+}19]. Multi-Modal [KLC15]. multi-threaded [VV00]. multicast [GL06]. multicriteria [DHLO09]. multidimensional [DM05]. Multilevel [HSS19b, RN11, SRB08, HSW08]. Multimodal [DPW15, DLL23]. Multipattern [STK06]. multiple [FN04, HFN05, KM13b, TRC11]. multiplication [ERS99].

naming [KM13a]. Nash [PS06]. national [BS99]. Near [BFKK20]. Near-Optimal [BFKK20]. Nearest [EP16a, EP16b]. Neighbor [EP16a, EP16b]. Network [GK01, HSS19b, QSL^{+}21]. Network-Flow-Based [GK01]. Networks [BBDW19, DLL23, EP16a, EP16b, GMS16, KLC15, LS15, dJvdLT23, ADGW13, BSWW04, EJ99, JMN99, LRAM06, Li08]. NN [SF18]. NN-Graph [SF18]. Node [BCD^{+}18, BSWW04]. Nodes [PGI18, NTB05]. Non [DLL23]. Non-scheduled [DLL23]. nonlinear [BLLW13]. number [HH02]. Numerical [RMH^{+}16, LRAM06].

O-Efficient [ATV01, HMP^{+}18, ADT03]. Oblivious [BM21, BFV08]. Obtaining [CKLM09]. off [KCC11]. off-the-shelf [KCC11]. offs [ACN23]. OLED [EKX09]. on-line [SWW00]. one [DT09, FMRT02]. Online [BKP20, AS02, ERW09, KM13b, Li08]. operations [Rad98]. optical [Li08]. Optimal [BFKK20, Hof13, PPM16, BMS09, CKLM09, EJ99, FN04]. Optimization [CDF^{+}23, HS18, BLLW13, DHLO09]. optimizing [LLH^{+}22]. Order [SF18, LOMSS05]. order-preserving [LOMSS05]. ordering [SRB08]. other [Epp00]. out-of-core [KCC11]. overlay [HSW08].

Packing [BLW16, BBJP08, LMMM05]. Paging [MNNW15]. Pair [NZ01]. Paired [MO15]. Pairs [CNV22, Epp00, VSM03]. Papers [DT08, MW09, SÁ09]. Parallel [BBO17, HT17, BHJ96, HJB98, KCC11]. parallelism [HFN05, NR00, VV00]. Parallelization [ABC^{+}18]. Parallelizing [IKM^{+}02]. Pareto [HS18]. Parsing [MRS01]. Partial [SFKM20]. Partition [CDF^{+}23, FKKS23]. Partitioning [ADGM06, FS22, GMS16, HNS20, HSS19b, MSS^{+}06, SSS15, SHG^{+}22, YBA^{+}22, CKM00, Kim99]. Path [DDF^{+}15, GO15, KKT^{+}20, BSWW04, CGG^{+}09, FPPR06, HTVW08, HSWW05, HSW08, MSM09, Shi00, WWZ05]. Path-decomposed [KKT^{+}20]. pathfinding [NAH04]. Paths [FR15, KLC15, ACN23, EJ02, VSM03]. Pathwidth [CMN16]. Pattern [LPR21, Kim99, KM13b]. PC [FPR23]. PC-Trees [FPR23]. Perfect [Bel23, BBPV11, DHW08]. Performance [BKS00, GK01, RVHE18, DM05, LL96, Li08, NAH04, XZK00]. Persistence [DSW19, KMN17]. persistent [ADT03]. personalized [BHJ96]. Phase [LMP23]. place [SW10]. placement [LFLSW08]. Planar [CHW19, FEMPS16a, FEMPS16b, PPM16, VH02, ADGM06, FHH^{+}00, HH08, HSW^{+}09, TMH11]. Planarity [CZ15]. Plane [AGG^{+}23, CDF^{+}23, FKKS23]. planes [Lev00]. planned [BS99]. Planning [DPW15, GRST12]. Plexes [BHM^{+}19]. Point [PPM16, ADT03, HH08, SSMJ99, VH02]. points [ECHS09]. Poisson [NG22]. Pólya [RMH^{+}16]. Polygonization [PPM16]. Polylogarithmic [IKRT01]. polynomial [CEM13]. polynomial-delay [CEM13]. Portable [LG02]. Potential [KBJ19]. Power [HRSZ98, MS02]. PQ [FPR23]. PQ-Trees [FPR23]. Practical [AN10, ACF^{+}21, BHJ96, CEF23, DEF^{+}21, GKN17, HNSS18, MNG16, CT09]. Practice [AGG^{+}23, BC19, CZ15, BW02, BBPV11, TW09]. precomputed [MSM09]. Predicting [Git96]. prediction [BNWG08]. Preface [AI08, Dem09, McG10, MW09, RS08, RM05, SÁ09, Vah11]. preference [Mic11]. prefix [BFK^{+}03]. Preparation [GC20]. Presence [DW15]. preserving [LOMSS05]. Price [BLW16]. primal [BLMM05]. primal-dual [BLMM05]. priority [BCFM00, CS00, GT04, San00, YZ98]. probabilistic [LRAM06]. Probability [NG22]. Problem [ABC^{+}18, BSNA21, CHW19, CAGDPS^{+}23, GJR22, PW17, BSTU08, BLMM05, BGL03, CEM13, CMP^{+}08, DL10, DS13, DK05, FT06, FMRT02, FPPR06, Jul09, LNW08, Pet03, SOS05, dSLAM05]. Problems [ATV01, KBJ19, FMRT02, LMS10, NPS^{+}11, SRB08]. Process [KST20]. processors [CT09]. Product [Hed15]. Programming [AKK19, LMP23, MW23, DS13, Hof13, LFLSW08, TPT^{+}11]. Projections [LMP23]. Propagation [LLH^{+}22]. properties [HH11]. Property [CILP20, Hed15]. protein [LFLSW08]. protocol [LRAM06]. prototyping [BKL00]. pseudo [ARMS08]. pseudo-Boolean [ARMS08]. psort [BBP11]. PTAS [TMH11]. public [PSWZ08, SWW00]. pure [PS06]. push [CGM^{+}98]. PVM [IKM^{+}02].

Quadratic [KST20]. Quality [SHG^{+}22]. Quantum [JLM^{+}21]. Quasi [NPR^{+}19]. Quasi-wideness [NPR^{+}19]. Quasirandom [DFKS11]. Queries [BDSU22, EP16a, EP16b, HSW08, MSM09]. queues [BCFM00, CS00, GT04, San00]. Quick [HHS22]. QUICKSORT} [ES02, BFM08, CT09, EW19]. quite [AW09].

Radix [RR01]. radixsort [AN98]. railroad [SWW00]. Random [BN19, LMP23, Man18, VM18, SZ05, TRC11]. Randomized [DW15, PPM16]. Ranking [CW09, vdBG^{+}22, LG02, VV00]. Rapid [BKL00]. reaction [KM13a]. Reactive [BP97]. Real [BSW19, ELS13, KKN^{+}21, FNP06, NPS^{+}11]. Real-time [BSW19]. Real-World [ELS13, KKN^{+}21, NPS^{+}11]. realistic [JMN99]. Reallocation [FCLMC21]. Rearrangement [GD15]. Recognition [CFP19, BG97]. reconfigurable [dSLAM05]. reconstruct [DFKM09]. reconstruction [HNR^{+}99]. reconstructions [PF08]. recovery [BA06]. Rectangle [HLL06]. Rectangles [SFKM20]. rectangular [LMMM05]. Rectilinear [RRRW19, YBA^{+}22]. Redesigning [AZ10]. Redistricting [LF19]. reduce [AN10]. Reduction [ECHS09, KKN^{+}21, Ull15, BW02, GGHN08].

Reference [HHS22]. Refinement [HSS19b, KCC11, NAH04]. refinement-based [NAH04]. Region [Man18]. Regions [EOP20]. registers [WACV02]. ReHub [EP16a, EP16b]. Relational [Ner02]. relinking [FPPR06]. reordering [ERW09]. Repetitive [NP16]. Representations [BBO17]. requirement [AN10]. resilient [FPGI13]. restrictions [GRST12]. Retrieval [GKN17, LMP23, NPV15, Ull15]. Reverse [BCF^{+}21, EP16a, EP16b]. Reverse-Safe [BCF^{+}21]. Revisited [Bel23, FPR09, HLL06, PU11]. right [Lev00]. Rips [DSW19]. Rips-filtration [DSW19]. Road [BBDW19, LS15, ADGW13]. Robust [DLY23, BD09]. Rotating [GNR16]. Rounding [DW15]. Route [DPW15, GRST12]. Routes [LS15, ADGW13]. Routing [BFKK20, DLL23, BD09, JMN99, Li08]. Rule [QSL^{+}21]. Rule-based [QSL^{+}21]. rules [BLLW13, ECHS09]. rumor [DFKS11]. Runtimes [LW22].

S [Mcg16]. Safe [BCF^{+}21]. salesman [FMRT02]. Sample [GC20]. Sampling [JB20, SZ05]. SAT [BP97, CZ15, SS23, TS16]. SAT-boosted [SS23]. satisfaction [Ull10]. Satisfiability [AON15, PW17, Spe15, IKM^{+}02, Spe10]. Scalable [HSS19a, vdBG^{+}22]. Scale [EP16a, EP16b]. Scan [DPSW18]. schedule [BS99]. scheduled [DLL23]. Scheduler [GK01]. Scheduling [FCLMC21, Lib01, AS02]. Schemes [SSS15]. SEA [D’A19, Fes13, GK16]. Search [HNS20, NG22, SS23, AW09, BP97, CW09, DM05, FCNP09, Git96, IKM^{+}02, NTB05, PU11, RN11, SOS05]. Searching [KM17, BLOLS09, Boy11, WEM11]. Section [MW09, SÁ09, Nik06]. seek [VH02]. Selected [MW09, SÁ09]. selection [Hof13]. selective [PF08]. Semidefinite [MW23]. sensitive [BP97]. sensor [LRAM06]. Separated [NZ01]. separator [HSW^{+}09]. Separators [FEMPS16a, FEMPS16b]. Sequences [BBG^{+}17, CILP20]. Sequential [JB20]. Set [BSNA21, GJR22, BLOLS09, FT06, SSMJ99]. set-covering [FT06]. Sets [BLN22, HGH21, HSS19a, LS15, MNG16, PPM16, KMY03, SZ04, VH02]. sgen1 [Spe10]. SHARC [BD09]. shelf [KCC11]. SHOP [FKKS23]. Short [FEMPS16a, FEMPS16b]. shortcutting [DT09]. Shortest [CGG^{+}09, DDF^{+}15, FR15, KLC15, ACN23, BSWW04, HSWW05, HSW08, MSM09, Shi00, VSM03, WWZ05]. Shortest-Path [DDF^{+}15, CGG^{+}09, BSWW04, HSWW05, HSW08, MSM09, WWZ05]. shortest-paths [VSM03]. sided [Mic11]. SimBa [DSW19]. Simple [FEMPS16a, FEMPS16b, FR15, ACN23, HH11, MPR04]. Simplicial [DSW19]. Simulated [CEF23]. Single [GC20, DK05, FN04, FINP98, Mic11]. Single-Droplet [GC20]. single-sided [Mic11]. single-source [DK05]. six [DL10]. skip [CS98]. SLS [LW22]. small [Spe10]. Smaller [GL20, GL22]. smoothness [GMS^{+}13]. software [BBP11, BKL00, KCC11, MHS99]. solution [AW09]. Solutions [TS16, CEM13]. solve [CMP^{+}08]. Solver [RVHE18]. Solvers [TS16]. Solving [FMRT02, KBJ19, NPS^{+}11]. Sort [RR01, PK06]. Sorting [NLG^{+}22, BHJ96, BBP11, BNWG08, BFV08, HJB98, Hof13, LOMSS05, MP08, RR00, SZ04, SZ05, SZR06, SW10, WACV02, XZK00]. source [DK05]. Space [BBO17, NPV15, ACN23, AN10, PSS09]. Space-Efficient [BBO17, PSS09]. spaces [FCNP09]. Spanners [AGG^{+}23, FG09]. Spanning [FMM20, NZ01, Jul09, WSdC00]. Sparse [ELS13, KST20]. spatial [FCNP09, HSS08, NR08]. Special [BM15, D’A19, GK16, Nik06, PC18, SZ16a, SZ16b, Zar19, Fes13, SÁ09]. Spectral [MW23]. speech [BG97]. speed [BDS^{+}10, Git96, HSWW05]. speed-up [BDS^{+}10, HSWW05]. Speeding [FCNP09, TPT^{+}11]. speedup [MSS^{+}06]. sphere [WEM11]. spreading [DFKS11]. squares [BBJP08]. SSLC [NG22]. Stable [BIS11, BBP11, IM09]. starting [AW09]. State [BBG^{+}17]. States [HRSZ98]. Static [vdBG^{+}22]. Statistical [AON15]. Steiner [TMH11, AAS^{+}19, BC19]. STL [FPR09]. Stochastic [HH11]. storage [GO13]. Strategies [GNR16, ERW09, SOS05]. stratified [PF08]. Streaming [FS22]. streams [HSS08]. String [CFP19, AZ10, FN04, Git96, HFN05, Kim99, LOMSS05, NR00, STK06, SZ05, SZR06, SW10]. string-pattern [Kim99]. Strings [HT17, LPR21, BFK^{+}03, SZ04]. strip [LMMM05]. Strong [BC19]. Structural [dJvdLT23]. Structure [LLH^{+}22, Rea12]. Structures [BDSU22, CS00, FPGI13, IGA05, LFLSW08, MS02, NG10]. Study [BLN22, BKP20, CZ15, FR15, FMZ01, IKRT01, Man18, PPM16, AS02, ACI97, BNWG08, BIS11, BKS00, BCFM00, CGM^{+}98, CMP^{+}08, ERS99, FG09, FPGI13, HH08, JMN99, Jac10, KZ08, LG02, LRAM06, Mag98, MPR04, MHS00, NAH04, PCJ97, SWW00, SSMJ99]. Subgraph [CHW19, Ull10]. Subgraphs [CDF^{+}23, FKKS23]. Subgroup [Hed15]. Subjects [SBG01]. Sublinear [BM19, VM18]. subsequence [BGL03]. substring [BSTU08]. subtree [Rad98]. Succinct [BBO17]. Suffix [BBO17, BFO16a, BFO16b, CILP20, GNF15, NP16, NLG^{+}22, DKMS08, GO13, MP08, NR00, VMGD09]. Suitor [ABM22]. Sum [BBJP08]. Sum-of-squares [BBJP08]. Summarizing [HSS08]. Supercomputer [BGQ19]. Support [SSS21]. supports [IJS^{+}06]. surfaces [HTVW08, LNW08]. switches [TRC11]. Symmetric [YBA^{+}22]. Symmetry [ARMS08]. synchronization [BA06]. Systems [Lib01, PSWZ08, dSLAM05].

table [AZ10, PF08]. Tabu [SS23, SOS05]. tailed [LW22]. takes [DFKM09]. teacher [SOS05]. technique [BLMM05]. techniques [BDS^{+}10, FT06, HSWW05, KM13a, NAH04]. Temporal [BHM^{+}19]. Terracost [HTVW08]. terrain [LNW08]. Terrains [ATV01, HTVW08]. test [FT06]. test-cover [FT06]. Testing [CZ15]. Text [BCF^{+}21, AN10, BLOLS09]. Theorem [RMH^{+}16]. Theory [BBPV11, FMM20, KKN^{+}21]. threaded [VV00]. three [SSMJ99]. three-dimensional [SSMJ99]. threshold [BSTU08]. throughput [Li08]. Tight [HGH21]. Time [BM19, KLC15, VM18, ACN23, BGSV13, BSW19]. Time-Dependent [KLC15, BGSV13]. times [BGSV13]. timetable [PSWZ08]. timetabling [SOS05]. Todinca [KBJ19]. Tool [DSW19, FGLP21, GD15]. Top [GKN17]. Top- [GKN17]. topological [PK06]. Topologies [BBO17]. Total [ABC^{+}18]. tournaments [CW09]. Trade [ACN23]. Trade-offs [ACN23]. Traffic [BSW19, BMS09]. Trajectories [CEF23]. Transitive [FMZ01, KZ08, PP06]. Translation [JM15]. transport [SWW00]. Transportation [DLL23, BSWW04, JMN99, PSWZ08]. travel [BGSV13]. traveling [FMRT02]. Treatment [SBG01]. Tree [BBO17, BC19, DDF^{+}15, DEF^{+}21, FMM20, GMS16, TMH11, DT09, EJ99, HNR^{+}99, Jul09, VMGD09]. Tree-Based [GMS16]. Trees [AAS^{+}19, BM21, NZ01, NP16, ADT03, CKLM09, CS98, EJ02, GO13, HW11, NR08, PF08, Rad98, TW09, WSdC00, FPR23]. Treewidth [ZP19]. trie [AZ10]. Tries [ACF^{+}21, GO15, KKT^{+}20, SZ04, SZ05]. Triple [Hed15]. Triple-Product-Property [Hed15]. Triples [Hed15]. Triplet [BM21]. true [SMEDM08]. TSP [DT09, Lev00, RVHE18]. Tuning [AKK19]. Two [CMTW19, KM13b, SMEDM08]. two-dimensional [KM13b]. Twol [GT04]. Twol-amalgamated [GT04].

UK [MO15]. Uncertainty [FMM20]. Understanding [LW22]. Undirected [BLW16]. Unexplored [EOP20]. unidirectional [BD09]. Uniform [CAGDPS^{+}23, NPR^{+}19, BLMM05]. unit [CGM^{+}98]. unit-capacity [CGM^{+}98]. United [HRSZ98]. unsplittable [DK05]. unweighted [EJ02]. Update [BM19]. Update-Time [BM19]. Updated [GHS21]. Updates [DDF^{+}15, GLLO22]. Updating [VM18]. Upgrading [Hed15]. Upward [CZ15, CGMW10]. User [DPW15]. User-Constrained [DPW15]. Using [BSW19, CMTW19, SF18, SZ05, ADT03, BKL00, CGM09, DM05, HSS08, IKM^{+}02, Kim99, KCC11, KM13b, KMY03, LOMSS05, MSM09, SZR06, TRC11, WACV02]. utilization [ERS99].

Value [TSP18]. values [GO13]. Variance [TSP18]. Vector [SSS21, Ull10]. vectorization [IKM^{+}02]. versatile [MP08]. versus [JLM^{+}21]. Vertex [CAGDPS^{+}23, DL10]. via [BN19, DSW19, Hof13, KBJ19, NZ01, PF08]. Virtual [JM15]. Visibility [ABC^{+}18]. Visibility-Index [ABC^{+}18]. VLSI [CKM00]. Volume [CEF23].

Wasserstein [DLY23]. Waste [GC20]. Waste-Efficient [GC20]. wavelength [EJ99, Li08]. Wavelet [DEF^{+}21]. WDM [Li08]. WEA [SÁ09]. Weak [CFP19]. Weakening [Spe15]. Weight [CS98, DT09]. Weight-biased [CS98]. Weighted [ABM22, CILP20, EJ02, FPPR06, LNW08, MS02, MHS99, MHS00, PS06]. weights [ADGM06]. Well [NZ01]. Well-Separated [NZ01]. wideness [NPR^{+}19]. width [NTB05, SSMJ99]. Windows [FCLMC21]. workload [BMS09]. World [ELS13, KKN^{+}21, NPS^{+}11].

XMT [VV00]. Xor [GL20, GL22].

Ziv [AN10, ACF^{+}21, KKP16a, KKP16b]. Ziv-based [AN10].

References

Ahmed:2019:MLS

[AAS^{+}19] Reyan Ahmed, Patrizio Angelini, Faryad Darabi Sahneh, Alon Efrat, David Glickenstein, Martin Gronemann, Niklas Heinsohn, Stephen G. Kobourov, Richard Spence, Joseph Watkins, and Alexander Wolff. Multi-level Steiner trees. ACM Journal of Experimental Algorithmics, 24(2):2.5:1–2.5:??, December 2019. CODEN ???? ISSN 1084-6654.

Auer:2015:EMC

[AB15] B. O. Fagginger Auer and R. H. Bisseling. Efficient matching for column intersection graphs. ACM Journal of Experimental Algorithmics, 19(??):1.3:1–1.3:??, February 2015. CODEN ???? ISSN 1084-6654.

Afshani:2018:EAT

[ABC^{+}18] Peyman Afshani, Mark De Berg, Henri Casanova, Ben Karsin, Colin Lambrechts, Nodari Sitchinava, and Constantinos Tsirogiannis. An efficient algorithm for the 1D total visibilityindex problem and its parallelization. ACM Journal of Experimental Algorithmics, 23(??):2.3:1–2.3:??, 2018. CODEN ???? ISSN 1084-6654.

Angriman:2022:BDS

[ABM22] Eugenio Angriman, Michał Boroń, and Henning Meyerhenke. A batch-dynamic suitor algorithm for approximating maximum weighted matching. ACM Journal of Experimental Algorithmics, 27(1):1.6:1–1.6:??, 2022. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3529228.

Arroyuelo:2021:EPL

[ACF^{+}21] Diego Arroyuelo, Rodrigo Cánovas, Johannes Fischer, Dominik Köppl, Marvin Löbel, Gonzalo Navarro, and Rajeev Raman. Engineering practical Lempel–Ziv tries. ACM Journal of Experimental Algorithmics, 26(1):14:1–14:47, December 2021. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/doi/

10.1145/3481638.

Alberts:1997:ESD

[ACI97] David Alberts, Giuseppe Cattaneo, and Giuseppe F. Italiano. An empirical study of dynamic graph algorithms. ACM Journal of Experimental Algorithmics, 2:5:1–5:??, ???? 1997. CODEN ???? ISSN 1084-6654.

Zoobi:2023:FKS

[ACN23] Ali Al Zoobi, David Coudert, and Nicolas Nisse. Finding the k shortest simple paths: Time and space trade-offs. ACM Journal of Experimental Algorithmics, 28:1.11:1–1.11:??, 2023. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3626567.

Aleksandrov:2006:PPG

[ADGM06] Lyudmil Aleksandrov, Hristo Djidjev, Hua Guo, and Anil Maheshwari. Partitioning planar graphs with costs and weights. ACM Journal of Experimental Algorithmics, 11:1.5:1–1.5:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Abraham:2013:ARR

[ADGW13] Ittai Abraham, Daniel Delling, Andrew V. Goldberg, and Renato F. Werneck. Alternative routes in road networks. ACM Journal of Experimental Algorithmics, 18(1):1.3:1–1.3:??, December 2013. CODEN ???? ISSN 1084-6654.

Arge:2003:EPL

[ADT03] Lars Arge, Andrew Danner, and Sha-Mayn Teh. I/O-efficient point location using persistent B-trees. ACM Journal of Experimental Algorithmics, 8:1.2:1–1.2:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Anderson:2023:BDP

[AGG^{+}23] Frederick Anderson, Anirban Ghosh, Matthew Graham, Lucas Mougeot, and David Wisnosky. Bounded-degree plane geometric spanners in practice. ACM Journal of Experimental Algorithmics, 28:1.1:1–1.1:??, 2023. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3582497.

Arge:2008:P

[AI08] Lars Arge and Giuseppe F. Italiano. Preface. ACM Journal of Experimental Algorithmics, 12:2.1:1–2.1:??, June 2008. CODEN ???? ISSN 1084-6654.

Altmanova:2019:ETF

[AKK19] Katerina Altmanová, Dusan Knop, and Martin Koutecký. Evaluating and tuning n-fold integer programming. ACM Journal of Experimental Algorithmics, 24(2):2.2:1–2.2:??, 2019. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/ft_

gateway.cfm?id=3330137.

Andersen:2020:GFE

[AM20] Jakob L. Andersen and Daniel Merkle. A generic framework for engineering graph canonization algorithms. ACM Journal of Experimental Algorithmics, 25(1):1–26, April 2020. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/abs/
10.1145/3356020.

Andersson:1998:IR

[AN98] Arne Andersson and Stefan Nilsson. Implementing radixsort. ACM Journal of Experimental Algorithmics, 3:7:1–7:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Arroyuelo:2010:PAR

[AN10] Diego Arroyuelo and Gonzalo Navarro. Practical approaches to reduce the space requirement of Lempel–Ziv-based compressed text indices. ACM Journal of Experimental Algorithmics, 15(1): 15:1–15:??, March 2010. CODEN ???? ISSN 1084-6654.

Anonymous:2008:EGC

[Ano08] Anonymous. Engineering graph clustering: Models and experimental evaluation. ACM Journal of Experimental Algorithmics, 12:1.1:1–1.1:??, June 2008. CODEN ???? ISSN 1084-6654.

Angione:2015:SMB

[AON15] Claudio Angione, Annalisa Occhipinti, and Giuseppe Nicosia. Satisfiability by Maxwell–Boltzmann and Bose–Einstein statistical distributions. ACM Journal of Experimental Algorithmics, 19(??):1.4:1–1.4:??, February 2015. CODEN ???? ISSN 1084-6654.

Aloul:2008:SBP

[ARMS08] Fadi A. Aloul, Arathi Ramani, Igor L. Markov, and Karem A. Sakallah. Symmetry breaking for pseudo-Boolean formulas. ACM Journal of Experimental Algorithmics, 12:1.3:1–1.3:??, June 2008. CODEN ???? ISSN 1084-6654.

Albers:2002:ESO

[AS02] Susanne Albers and Bianca Schröder. An experimental study of online scheduling algorithms. ACM Journal of Experimental Algorithmics, 7:3, ???? 2002. CODEN ???? ISSN 1084-6654.

Arge:2001:EAP

[ATV01] Lars Arge, Laura Toma, and Jeffrey Scott Vitter. I/O-efficient algorithms for problems on grid-based terrains. ACM Journal of Experimental Algorithmics, 6:1:1–1:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Avdil:2009:LSS

[AW09] Alaubek Avdil and Karsten Weihe. Local search starting from an LP solution: Fast and quite good. ACM Journal of Experimental Algorithmics, 14(1):6:1–6:??, December 2009. CODEN ???? ISSN 1084-6654.

Askitis:2010:RSH

[AZ10] Nikolas Askitis and Justin Zobel. Redesigning the string hash table, burst trie, and BST to exploit cache. ACM Journal of Experimental Algorithmics, 15(1):17:1–17:??, March 2010. CO-DEN ???? ISSN 1084-6654.

Boukerche:2006:ICC

[BA06] Azzedine Boukerche and Alba Cristina Magalhaes Alves De Melo. Integrating coordinated checkpointing and recovery mechanisms into DSM synchronization barriers. ACM Journal of Experimental Algorithmics, 11:2.9:1–2.9:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Baum:2019:FEC

[BBDW19] Moritz Baum, Valentin Buchhold, Julian Dibbelt, and Dorothea Wagner. Fast exact computation of isocontours in road networks. ACM Journal of Experimental Algorithmics, 24(1): 1.18:1–1.18:??, October 2019. CODEN ???? ISSN 1084-6654.

Buchin:2017:CFD

[BBG^{+}17] Kevin Buchin, Maike Buchin, Joachim Gudmundsson, Michael Horton, and Stef Sijben. Compact flow diagrams for state sequences. ACM Journal of Experimental Algorithmics, 22(??): 1.7:1–1.7:??, 2017. CODEN ???? ISSN 1084-6654.

Bender:2008:SSH

[BBJP08] Michael A. Bender, Bryan Bradley, Geetha Jagannathan, and Krishnan Pillaipakkamnatt. Sum-of-squares heuristics for bin packing and memory allocation. ACM Journal of Experimental Algorithmics, 12:2.3:1–2.3:??, June 2008. CODEN ???? ISSN 1084-6654.

Baier:2017:SEP

[BBO17] Uwe Baier, Timo Beller, and Enno Ohlebusch. Space-efficient parallel construction of succinct representations of suffix tree topologies. ACM Journal of Experimental Algorithmics, 22(??): 1.1:1–1.1:??, 2017. CODEN ???? ISSN 1084-6654.

Bertasi:2011:PYA

[BBP11] Paolo Bertasi, Marco Bressan, and Enoch Peserico. psort, yet another fast stable sorting software. ACM Journal of Experimental Algorithmics, 16(1):24:1–24:??, 2011. CODEN ???? ISSN 1084-6654.

Belazzougui:2011:TPM

[BBPV11] Djamal Belazzougui, Paolo Boldi, Rasmus Pagh, and Sebastiano Vigna. Theory and practice of monotone minimal perfect hashing. ACM Journal of Experimental Algorithmics, 16(1): 32:1–32:??, 2011. CODEN ???? ISSN 1084-6654.

Beyer:2019:SST

[BC19] Stephan Beyer and Markus Chimani. Strong Steiner tree approximations in practice. ACM Journal of Experimental Algorithmics, 24(1):1.7:1–1.7:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3299903.

Bergamini:2018:IBC

[BCD^{+}18] Elisabetta Bergamini, Pierluigi Crescenzi, Gianlorenzo D’Angelo, Henning Meyerhenke, Lorenzo Severini, and Yllka Velaj. Improving the betweenness centrality of a node by adding links. ACM Journal of Experimental Algorithmics, 23(??):1.5:1– 1.5:??, 2018. CODEN ???? ISSN 1084-6654.

Bernardini:2021:RST

[BCF^{+}21] Giulia Bernardini, Huiping Chen, Gabriele Fici, Grigorios Loukides, and Solon P. Pissis. Reverse-safe text indexing. ACM Journal of Experimental Algorithmics, 26(1):1–26, December 2021. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/doi/10.1145/3461698.

Brengel:2000:ESP

[BCFM00] Klaus Brengel, Andreas Crauser, Paolo Ferragina, and Ulrich Meyer. An experimental study of priority queues in external memory. ACM Journal of Experimental Algorithmics, 5:17:1– 17:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Bauer:2009:SFR

[BD09] Reinhard Bauer and Daniel Delling. SHARC: Fast and robust unidirectional routing. ACM Journal of Experimental Algorithmics, 14:4:1–4:??, May 2009. CODEN ???? ISSN 1084-6654.

Bauer:2010:CHG

[BDS^{+}10] Reinhard Bauer, Daniel Delling, Peter Sanders, Dennis Schieferdecker, Dominik Schultes, and Dorothea Wagner. Combining hierarchical and goal-directed speed-up techniques for Dijkstra’s algorithm. ACM Journal of Experimental Algorithmics, 15(1): 2.3:1–2.3:??, March 2010. CODEN ???? ISSN 1084-6654.

Bertrand:2022:ADS

[BDSU22] Jules Bertrand, Fanny Dufossé, Somesh Singh, and Bora Uçar. Algorithms and data structures for hyperedge queries. ACM Journal of Experimental Algorithmics, 27(1):1.13:1–1.13:??, 2022. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/doi/10.1145/3568421.

Beling:2023:FBM

[Bel23] Piotr Beling. Fingerprinting-based minimal perfect hashing revisited. ACM Journal of Experimental Algorithmics, 28:1.4:1– 1.4:??, 2023. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/doi/10.1145/3596453.

Buchsbaum:2003:FPM

[BFK^{+}03] Adam L. Buchsbaum, Glenn S. Fowler, Balachannder Kirishnamurthy, Kiem-Phong Vo, and Jia Wang. Fast prefix matching of bounded strings. ACM Journal of Experimental Algorithmics, 8:1.3:1–1.3:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Blasius:2020:HEN

[BFKK20] Thomas Bläsius, Tobias Friedrich, Maximilian Katzmann, and Anton Krohmer. Hyperbolic embeddings for near-optimal greedy routing. ACM Journal of Experimental Algorithmics, 25(1):1–18, April 2020. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/abs/10.1145/3381751.

Brodal:2008:AQ

[BFM08] Gerth Stølting Brodal, Rolf Fagerberg, and Gabriel Moruz. On the adaptiveness of Quicksort. ACM Journal of Experimental Algorithmics, 12:3.2:1–3.2:??, June 2008. CODEN ???? ISSN 1084-6654.

Bingmann:2016:ISL

[BFO16a] Timo Bingmann, Johannes Fischer, and Vitaly Osipov. Inducing suffix and LCP arrays in external memory. ACM Journal of Experimental Algorithmics, 21(1):2.3:1–2.3:??, November 2016. CODEN ???? ISSN 1084-6654.

Bingmann:2017:ISL

[BFO16b] Timo Bingmann, Johannes Fischer, and Vitaly Osipov. Inducing suffix and LCP arrays in external memory. ACM Journal of Experimental Algorithmics, 21(1):2.3:1–2.3:??, 2016. CODEN ???? ISSN 1084-6654.

Brodal:2008:ECO

[BFV08] Gerth Stølting Brodal, Rolf Fagerberg, and Kristoffer Vinther. Engineering a cache-oblivious sorting algorithm. ACM Journal of Experimental Algorithmics, 12:2.2:1–2.2:??, June 2008. CO-DEN ???? ISSN 1084-6654.

Buchsbaum:1997:AAS

[BG97] Adam L. Buchsbaum and Raffaele Giancarlo. Algorithmic aspects in speech recognition: an introduction. ACM Journal of Experimental Algorithmics, 2:1:1–1:??, ???? 1997. CODEN ???? ISSN 1084-6654.

Breimer:2003:LAL

[BGL03] Eric A. Breimer, Mark K. Goldberg, and Darren T. Lim. A learning algorithm for the longest common subsequence problem. ACM Journal of Experimental Algorithmics, 8:2.1:1–2.1:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Bjorklund:2019:FHF

[BGQ19] Andreas Björklund, Brajesh Gupt, and Nicolás Quesada. A faster Hafnian formula for complex matrices and its benchmarking on a supercomputer. ACM Journal of Experimental Algorithmics, 24(1):1.11:1–1.11:??, October 2019. CODEN ????

ISSN 1084-6654. URL https://dl.acm.org/ft_gateway.cfm?
id=3325111.

Batz:2013:MTD

[BGSV13] G. Veit Batz, Robert Geisberger, Peter Sanders, and Christian Vetter. Minimum time-dependent travel times with contraction hierarchies. ACM Journal of Experimental Algorithmics, 18(1): 1.4:1–1.4:??, December 2013. CODEN ???? ISSN 1084-6654.

Bader:1996:PPA

[BHJ96] David A. Bader, David R. Helman, and Joseph JáJá. Practical parallel algorithms for personalized communication and integer sorting. ACM Journal of Experimental Algorithmics, 1:3:1–3:??, ???? 1996. CODEN ???? ISSN 1084-6654.

Bentert:2019:LAM

[BHM^{+}19] Matthias Bentert, Anne-Sophie Himmel, Hendrik Molter, Marco Morik, Rolf Niedermeier, and René Saitenmacher. Listing all maximal k-plexes in temporal graphs. ACM Journal of Experimental Algorithmics, 24(1):1.13:1–1.13:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://dl.acm.org/

ft_gateway.cfm?id=3325859.

Biro:2011:SMC

[BIS11] Péter Biró, Robert W. Irving, and Ildikó Schlotter. Stable matching with couples: an empirical study. ACM Journal of Experimental Algorithmics, 16(1):12:1–12:??, 2011. CODEN ???? ISSN 1084-6654.

Boghossian:2000:RSP

[BKL00] N. P. Boghossian, O. Kohlbacher, and H. P. Lenhof. Rapid software prototyping in molecular modeling using the biochemical algorithms library (BALL). ACM Journal of Experimental Algorithmics, 5:16:1–16:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Borodin:2020:ESA

[BKP20] Allan Borodin, Christodoulos Karavasilis, and Denis Pankratov. An experimental study of algorithms for online bipartite matching. ACM Journal of Experimental Algorithmics, 25(1): 1–37, April 2020. CODEN ???? ISSN 1084-6654. URL https:

//dl.acm.org/doi/abs/10.1145/3379552.

Bojesen:2000:PEC

[BKS00] Jesper Bojesen, Jyrki Katajainen, and Maz Spork. Performance engineering case study: heap construction. ACM Journal of Experimental Algorithmics, 5:15:1–15:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Bonami:2013:BRC

[BLLW13] Pierre Bonami, Jon Lee, Sven Leyffer, and Andreas Wächter. On branching rules for convex mixed-integer nonlinear optimization. ACM Journal of Experimental Algorithmics, 18(1):2.6:1– 2.6:??, December 2013. CODEN ???? ISSN 1084-6654.

Bracht:2005:GAA

[BLMM05] Evandro C. Bracht, Luis, A. A. Meira, and F. K. Miyazawa. A greedy approximation algorithm for the uniform metric labeling problem analyzed by a primal-dual technique. ACM Journal of Experimental Algorithmics, 10:2.11:1–2.11:??, ???? 2005. CO-DEN ???? ISSN 1084-6654.

Bhore:2022:ASF

[BLN22] Sujoy Bhore, Guangping Li, and Martin Nöllenburg. An algorithmic study of fully dynamic independent sets for map labeling. ACM Journal of Experimental Algorithmics, 27(1):1.8:1– 1.8:??, 2022. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/doi/10.1145/3514240.

Barbay:2009:EIS

[BLOLS09] Jérémy Barbay, Alejandro López-Ortiz, Tyler Lu, and Alejandro Salinger. An experimental investigation of set intersection algorithms for text searching. ACM Journal of Experimental Algorithmics, 14(1):7:1–7:??, December 2009. CODEN ???? ISSN 1084-6654.

Bergner:2016:BPC

[BLW16] Martin Bergner, Marco E. Lübbecke, and Jonas T. Witt. A branch-price-and-cut algorithm for packing cuts in undirected graphs. ACM Journal of Experimental Algorithmics, 21(1): 1.2:1–1.2:??, November 2016. CODEN ???? ISSN 1084-6654.

Bader:2015:ISI

[BM15] David A. Bader and Petra Mutzel. Introduction to special issue ALENEX’12. ACM Journal of Experimental Algorithmics, 19 (??):3.1:1–3.1:??, February 2015. CODEN ???? ISSN 1084-6654.

Barenboim:2019:FDG

[BM19] Leonid Barenboim and Tzalik Maimon. Fully dynamic graph algorithms inspired by distributed computing: Deterministic maximal matching and edge coloring in sublinear updatetime. ACM Journal of Experimental Algorithmics, 24(1):1.14:1– 1.14:??, October 2019. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/ft_gateway.cfm?id=3338529.

Brodal:2021:COA

[BM21] Gerth Stølting Brodal and Konstantinos Mampentzidis. Cache oblivious algorithms for computing the triplet distance between trees. ACM Journal of Experimental Algorithmics, 26(1):1–44, December 2021. CODEN ???? ISSN 1084-6654. URL https:/

/dl.acm.org/doi/10.1145/3433651.

Basu:2009:GAO

[BMS09] Amitabh Basu, Joseph S. B. Mitchell, and Girish Kumar Sabhnani. Geometric algorithms for optimal airspace design and air traffic controller workload balancing. ACM Journal of Experimental Algorithmics, 14:3:1–3:??, May 2009. CODEN ???? ISSN 1084-6654.

Borassi:2019:KAA

[BN19] Michele Borassi and Emanuele Natale. KADABRA is an ADaptive Algorithm for Betweenness via Random Approximation. ACM Journal of Experimental Algorithmics, 24(1):1.2:1–1.2:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/ft_gateway.cfm?id=3284359.

Biggar:2008:ESS

[BNWG08] Paul Biggar, Nicholas Nash, Kevin Williams, and David Gregg. An experimental study of sorting and branch prediction. ACM Journal of Experimental Algorithmics, 12:1.8:1–1.8:??, June 2008. CODEN ???? ISSN 1084-6654.

Boytsov:2011:IMA

[Boy11] Leonid Boytsov. Indexing methods for approximate dictionary searching: Comparative analysis. ACM Journal of Experimental Algorithmics, 16(1):1.1:1–1.1:??, May 2011. CODEN ???? ISSN 1084-6654.

Battiti:1997:RSH

[BP97] Roberto Battiti and Marco Protasi. Reactive search, a historysensitive heuristic for MAX-SAT. ACM Journal of Experimental Algorithmics, 2:2:1–2:??, ???? 1997. CODEN ???? ISSN 1084-6654.

Burke:1999:MAS

[BS99] E. K. Burke and A. J. Smith. A memetic algorithm to schedule planned maintenance for the national grid. ACM Journal of Experimental Algorithmics, 4:1:1–1:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Baharev:2021:EMM

[BSNA21] Ali Baharev, Hermann Schichl, Arnold Neumaier, and Tobias Achterberg. An exact method for the minimum feedback arc set problem. ACM Journal of Experimental Algorithmics, 26 (1):1–28, December 2021. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3446429. See com-

ments [GJR22].

Barsky:2008:GAT

[BSTU08] Marina Barsky, Ulrike Stege, Alex Thomo, and Chris Upton. A graph approach to the threshold all-against-all substring matching problem. ACM Journal of Experimental Algorithmics, 12: 1.10:1–1.10:??, June 2008. CODEN ???? ISSN 1084-6654.

Buchhold:2019:RTT

[BSW19] Valentin Buchhold, Peter Sanders, and Dorothea Wagner. Realtime traffic assignment using engineered customizable contraction hierarchies. ACM Journal of Experimental Algorithmics, 24(2):2.4:1–2.4:??, December 2019. CODEN ???? ISSN 1084-6654.

Brandes:2004:GNC

[BSWW04] Ulrik Brandes, Frank Schulz, Dorothea Wagner, and Thomas Willhalm. Generating node coordinates for shortest-path computations in transportation networks. ACM Journal of Experimental Algorithmics, 9:1.1:1–1.1:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Backes:2002:HLB

[BW02] Werner Backes and Susanne Wetzel. Heuristics on lattice basis reduction in practice. ACM Journal of Experimental Algorithmics, 7:1, ???? 2002. CODEN ???? ISSN 1084-6654.

Cornejo-Acosta:2023:CHU

[CAGDPS^{+}23] José Alejandro Cornejo-Acosta, Jesús García-Díaz, Julio César Pérez-Sansalvador, Roger Z. Ríos-Mercado, and Saúl Eduardo Pomares-Hernández. A constructive heuristic for the uniform capacitated vertex k-center problem. ACM Journal of Experimental Algorithmics, 28:1.6:1–1.6:??, 2023. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/doi/10.1145/3604911.

Cohen:2015:CGH

[CCL15] Nathann Cohen, David Coudert, and Aurélien Lancin. On computing the Gromov hyperbolicity. ACM Journal of Experimental Algorithmics, 20(??):1.6:1–1.6:??, 2015. CODEN ???? ISSN 1084-6654.

Crombez:2023:COB

[CDF^{+}23] Loïc Crombez, Guilherme D. Da Fonseca, Florian Fontan, Yan Gerard, Aldo Gonzalez-Lorenzo, Pascal Lafourcade, Luc Libralesso, Benjamin Momège, Jack Spalding-Jamieson, Brandon Zhang, and Da Wei Zheng. Conflict optimization for binary CSP applied to minimum partition into plane subgraphs and graph coloring. ACM Journal of Experimental Algorithmics, 28:1.2:1–1.2:??, 2023. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3588869.

Chalkis:2023:PAV

[CEF23] Apostolos Chalkis, Ioannis Z. Emiris, and Vissarion Fisikopoulos. A practical algorithm for volume estimation based on billiard trajectories and simulated annealing. ACM Journal of Experimental Algorithmics, 28:1.3:1–1.3:??, 2023. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3584182.

Canzar:2013:PDA

[CEM13] Stefan Canzar, Khaled Elbassioni, and Julián Mestre. A polynomial-delay algorithm for enumerating approximate solutions to the interval constrained coloring problem. ACM Journal of Experimental Algorithmics, 18(1):2.2:1–2.2:??, December 2013. CODEN ???? ISSN 1084-6654.

Cantone:2019:LES

[CFP19] Domenico Cantone, Simone Faro, and Arianna Pavone. Linear and efficient string matching algorithms based on weak factor recognition. ACM Journal of Experimental Algorithmics, 24(1):

1.8:1–1.8:??, October 2019. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/ft_gateway.cfm?id=3301295.

Cherkassky:2009:SPF

[CGG^{+}09] Boris V. Cherkassky, Loukas Georgiadis, Andrew V. Goldberg, Robert E. Tarjan, and Renato F. Werneck. Shortest-path feasibility algorithms: An experimental evaluation. ACM Journal of Experimental Algorithmics, 14:7:1–7:??, May 2009. CODEN ???? ISSN 1084-6654.

Cherkassky:1998:APC

[CGM^{+}98] Boris V. Cherkassky, Andrew V. Goldberg, Paul Martin, Joao C. Setubal, and Jorge Stolfi. Augment or push: a computational study of bipartite matching and unit-capacity flow algorithms. ACM Journal of Experimental Algorithmics, 3:8:1– 8:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Chimani:2009:EEC

[CGM09] Markus Chimani, Carsten Gutwenger, and Petra Mutzel. Experiments on exact crossing minimization using column generation. ACM Journal of Experimental Algorithmics, 14(1):4:1– 4:??, December 2009. CODEN ???? ISSN 1084-6654.

Chimani:2010:LFU

[CGMW10] Markus Chimani, Carsten Gutwenger, Petra Mutzel, and Hoi-Ming Wong. Layer-free upward crossing minimization. ACM Journal of Experimental Algorithmics, 15(1):2.2:1–2.2:??, March 2010. CODEN ???? ISSN 1084-6654.

Chimani:2019:EAM

[CHW19] Markus Chimani, Ivo Hedtke, and Tilo Wiedera. Exact algorithms for the maximum planar subgraph problem: New models and experiments. ACM Journal of Experimental Algorithmics, 24(1):2.1:1–2.1:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3320344.

Charalampopoulos:2020:PSA

[CILP20] Panagiotis Charalampopoulos, Costas S. Iliopoulos, Chang Liu, and Solon P. Pissis. Property suffix array with applications in indexing weighted sequences. ACM Journal of Experimental Algorithmics, 25(1):1–16, April 2020. CODEN ???? ISSN 1084-

6654. URL https://dl.acm.org/doi/10.1145/3385898.

Chimani:2009:OOC

[CKLM09] Markus Chimani, Maria Kandyba, Ivana Ljubić, and Petra Mutzel. Obtaining optimal k-cardinality trees fast. ACM Journal of Experimental Algorithmics, 14(1):2.5:1–2.5:23, December 2009. CODEN ???? ISSN 1084-6654.

Caldwell:2000:DIM

[CKM00] Andrew E. Caldwell, Andrew B. Kahng, and Igor L. Markov. Design and implementation of move-based heuristics for VLSI hypergraph partitioning. ACM Journal of Experimental Algorithmics, 5:5:1–5:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Coudert:2016:EEB

[CMN16] David Coudert, Dorian Mazauric, and Nicolas Nisse. Experimental evaluation of a branch-and-bound algorithm for computing pathwidth and directed pathwidth. ACM Journal of Experimental Algorithmics, 21(1):1.3:1–1.3:??, November 2016. CODEN ???? ISSN 1084-6654.

Codenotti:2008:ESD

[CMP^{+}08] Bruno Codenotti, Benton Mccune, Sriram Pemmaraju, Rajiv Raman, and Kasturi Varadarajan. An experimental study of different approaches to solve the market equilibrium problem. ACM Journal of Experimental Algorithmics, 12:3.3:1–3.3:??, June 2008. CODEN ???? ISSN 1084-6654.

Cazals:2019:CTC

[CMTW19] F. Cazals, D. Mazauric, R. Tetley, and R. Watrigant. Comparing two clusterings using matchings between clusters of clusters. ACM Journal of Experimental Algorithmics, 24(1):1.17:1– 1.17:??, October 2019. CODEN ???? ISSN 1084-6654.

Coudert:2022:EFA

[CNV22] David Coudert, André Nusser, and Laurent Viennot. Enumeration of far-apart pairs by decreasing distance for faster hyperbolicity computation. ACM Journal of Experimental Algorithmics, 27(1):1.15:1–1.15:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3569169.

Cho:1998:WBL

[CS98] Seonghun Cho and Sartaj Sahni. Weight-biased leftist trees and modified skip lists. ACM Journal of Experimental Algorithmics, 3:2:1–2:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Chong:2000:CBD

[CS00] Kyn-Rak Chong and Sartaj Sahni. Correspondence-based data structures for double-ended priority queues. ACM Journal of Experimental Algorithmics, 5:2:1–2:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Cederman:2009:GQP

[CT09] Daniel Cederman and Philippas Tsigas. GPU-Quicksort: a practical Quicksort algorithm for graphics processors. ACM Journal of Experimental Algorithmics, 14(1):4:1–4:??, December 2009. CODEN ???? ISSN 1084-6654.

Coleman:2009:RTL

[CW09] Tom Coleman and Anthony Wirth. Ranking tournaments: Local search and a new algorithm. ACM Journal of Experimental Algorithmics, 14:6:1–6:??, May 2009. CODEN ???? ISSN 1084-6654.

Chimani:2015:UPT

[CZ15] Markus Chimani and Robert Zeranski. Upward planarity testing in practice: SAT formulations and comparative study. ACM Journal of Experimental Algorithmics, 20(??):1.2:1–1.2:??, 2015. CODEN ???? ISSN 1084-6654.

DAngelo:2019:ISI

[D’A19] Gianlorenzo D’Angelo. Introduction to the special issue SEA 2018. ACM Journal of Experimental Algorithmics, 24(2):2.1:1– 2.1:??, December 2019. CODEN ???? ISSN 1084-6654.

DAndrea:2015:DMS

[DDF^{+}15] Annalisa D’Andrea, Mattia D’Emidio, Daniele Frigioni, Stefano Leucci, and Guido Proietti. Dynamic maintenance of a shortestpath tree on homogeneous batches of updates: New algorithms and experiments. ACM Journal of Experimental Algorithmics, 20(??):1.5:1–1.5:??, 2015. CODEN ???? ISSN 1084-6654.

DAngelo:2019:FDH

[DDF19] Gianlorenzo D’Angelo, Mattia D’Emidio, and Daniele Frigioni. Fully dynamic 2-hop cover labeling. ACM Journal of Experimental Algorithmics, 24(1):1.6:1–1.6:??, October 2019. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/ft_

gateway.cfm?id=3299901.

Dinklage:2021:PWT

[DEF^{+}21] Patrick Dinklage, Jonas Ellert, Johannes Fischer, Florian Kurpicz, and Marvin Löbel. Practical wavelet tree construction. ACM Journal of Experimental Algorithmics, 26(1):1–67, December 2021. CODEN ???? ISSN 1084-6654. URL https:/

/dl.acm.org/doi/10.1145/3457197.

Demetrescu:2009:P

[Dem09] Camil Demetrescu. Preface. ACM Journal of Experimental Algorithmics, 14(1):1:1–1:??, December 2009. CODEN ???? ISSN 1084-6654.

Demestrescu:2001:BCM

[DF01] Camil Demestrescu and Irene Finocchi. Breaking cycles for minimizing crossings. ACM Journal of Experimental Algorithmics, 6:2:1–2:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Dumitriu:2009:HMG

[DFKM09] Daniel Dumitriu, Stefan Funke, Martin Kutz, and Nikola Milosavljević. How much geometry it takes to reconstruct a 2-manifold in R^{3}. ACM Journal of Experimental Algorithmics, 14:2:1–2:??, May 2009. CODEN ???? ISSN 1084-6654.

Doerr:2011:QRS

[DFKS11] Benjamin Doerr, Tobias Friedrich, Marvin Künnemann, and Thomas Sauerwald. Quasirandom rumor spreading: an experimental analysis. ACM Journal of Experimental Algorithmics, 16(1):33:1–33:??, 2011. CODEN ???? ISSN 1084-6654.

DeLoera:2009:CMM

[DHLO09] Jesús A. De Loera, David C. Haws, Jon Lee, and Allison O’Hair. Computation in multicriteria matroid optimization. ACM Journal of Experimental Algorithmics, 14(1):8:1–8:??, December 2009. CODEN ???? ISSN 1084-6654.

Dietzfelbinger:2008:DIB

[DHW08] Martin Dietzfelbinger, Martin Hühne, and Christoph Weidling. A dictionary implementation based on dynamic perfect hashing. ACM Journal of Experimental Algorithmics, 12:1.11:1–1.11:??, June 2008. CODEN ???? ISSN 1084-6654.

deJong:2023:AEC

[dJvdLT23] Rachel G. de Jong, Mark P. J. van der Loo, and Frank W. Takes. Algorithms for efficiently computing structural anonymity in complex networks. ACM Journal of Experimental Algorithmics, 28:1.7:1–1.7:??, 2023. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3604908.

Du:2005:IAA

[DK05] Jingde Du and Stavros G. Kolliopoulos. Implementing approximation algorithms for the single-source unsplittable flow problem. ACM Journal of Experimental Algorithmics, 10:2.3:1– 2.3:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Dementiev:2008:BEM

[DKMS08] Roman Dementiev, Juha Kärkkäinen, Jens Mehnert, and Peter Sanders. Better external memory suffix array construction. ACM Journal of Experimental Algorithmics, 12:3.4:1– 3.4:??, June 2008. CODEN ???? ISSN 1084-6654.

Delbot:2010:AEC

[DL10] François Delbot and Christian Laforest. Analytical and experimental comparison of six algorithms for the vertex cover problem. ACM Journal of Experimental Algorithmics, 15(1):14:1– 14:??, March 2010. CODEN ???? ISSN 1084-6654.

Drakulic:2023:RMT

[DLL23] Darko Drakulic, Christelle Loiodice, and Vassilissa Lehoux. Routing in multimodal transportation networks with nonscheduled lines. ACM Journal of Experimental Algorithmics, 28:2.3:1–2.3:??, 2023. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3632969.

Ding:2023:DDA

[DLY23] Hu Ding, Wenjie Liu, and Mingquan Ye. A data-dependent approach for high-dimensional (robust) Wasserstein alignment. ACM Journal of Experimental Algorithmics, 28:1.8:1–1.8:??, 2023. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/doi/10.1145/3604910.

Duch:2005:IPM

[DM05] Amalia Duch and Conrado Martínez. Improving the performance of multidimensional search using fingers. ACM Journal of Experimental Algorithmics, 10:2.4:1–2.4:??, ???? 2005. CO-DEN ???? ISSN 1084-6654.

Dibbelt:2018:CSA

[DPSW18] Julian Dibbelt, Thomas Pajor, Ben Strasser, and Dorothea Wagner. Connection scan algorithm. ACM Journal of Experimental Algorithmics, 23(??):1.7:1–1.7:??, 2018. CODEN ???? ISSN 1084-6654.

Dibbelt:2015:UCM

[DPW15] Julian Dibbelt, Thomas Pajor, and Dorothea Wagner. Userconstrained multimodal route planning. ACM Journal of Experimental Algorithmics, 19(??):3.2:1–3.2:??, February 2015. CO-DEN ???? ISSN 1084-6654.

Delort:2013:HDP

[DS13] Charles Delort and Olivier Spanjaard. A hybrid dynamic programming approach to the biobjective binary knapsack problem. ACM Journal of Experimental Algorithmics, 18(1):1.2:1– 1.2:??, December 2013. CODEN ???? ISSN 1084-6654.

deSouza:2005:DMP

[dSLAM05] Cid C. de Souza, Andre M. Lima, Guido Araujo, and Nahri B. Moreano. The datapath merging problem in reconfigurable systems: Complexity, dual bounds and heuristic evaluation. ACM Journal of Experimental Algorithmics, 10:2.2:1–2.2:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Dibbelt:2016:CCH

[DSW16] Julian Dibbelt, Ben Strasser, and Dorothea Wagner. Customizable contraction hierarchies. ACM Journal of Experimental Algorithmics, 21(1):1.5:1–1.5:??, November 2016. CODEN ???? ISSN 1084-6654.

Dey:2019:SET

[DSW19] Tamal K. Dey, Dayu Shi, and Yusu Wang. SimBa: an efficient tool for approximating rips-filtration persistence via simplicial batch collapse. ACM Journal of Experimental Algorithmics, 24 (1):1.5:1–1.5:??, October 2019. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/ft_gateway.cfm?id=3284360.

Demetrescu:2008:PA

[DT08] Camil Demetrescu and Roberto Tamassia. Papers from ALENEX 2005. ACM Journal of Experimental Algorithmics, 12:3.1:1–3.1:??, June 2008. CODEN ???? ISSN 1084-6654.

Deineko:2009:FMW

[DT09] Vladimir Deineko and Alexander Tiskin. Fast minimum-weight double-tree shortcutting for metric TSP: Is the best one good enough? ACM Journal of Experimental Algorithmics, 14(1): 6:1–6:??, December 2009. CODEN ???? ISSN 1084-6654.

Doerr:2015:RRP

[DW15] Benjamin Doerr and Magnus Wahlström. Randomized rounding in the presence of a cardinality constraint. ACM Journal of Experimental Algorithmics, 19(??):1.2:1–1.2:??, February 2015. CODEN ???? ISSN 1084-6654.

Estivill-Castro:2009:RRD

[ECHS09] Vladimir Estivill-Castro, Apichat Heednacram, and Francis Suraweera. Reduction rules deliver efficient FPT-algorithms for covering points with lines. ACM Journal of Experimental Algorithmics, 14(1):7:1–7:??, December 2009. CODEN ???? ISSN 1084-6654.

Erlebach:1999:EIO

[EJ99] T. Erlebach and K. Jansen. Efficient implementation of an optimal greedy algorithm for wavelength assignment in directed tree networks. ACM Journal of Experimental Algorithmics, 4: 4:1–4:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Erlebach:2002:IAA

[EJ02] Thomas Erlebach and Klaus Jansen. Implementation of approximation algorithms for weighted and unweighted edge-disjoint paths in bidirected trees. ACM Journal of Experimental Algorithmics, 7:6, ???? 2002. CODEN ???? ISSN 1084-6654.

Eisenbrand:2009:ALO

[EKX09] Friedrich Eisenbrand, Andreas Karrenbauer, and Chihao Xu. Algorithms for longer OLED lifetime. ACM Journal of Experimental Algorithmics, 14(1):3:1–3:??, December 2009. CODEN ???? ISSN 1084-6654.

Eppstein:2013:LAM

[ELS13] David Eppstein, Maarten Löffler, and Darren Strash. Listing all maximal cliques in large sparse real-world graphs. ACM Journal of Experimental Algorithmics, 18(??):3.1:1–3.1:??, December 2013. CODEN ???? ISSN 1084-6654.

Escamocher:2020:GDC

[EOP20] Guillaume Escamocher, Barry O’Sullivan, and Steven David Prestwich. Generating difficult CNF instances in unexplored constrainedness regions. ACM Journal of Experimental Algorithmics, 25(1):1–12, April 2020. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3385651.

Efentakis:2016:REH

[EP16a] Alexandros Efentakis and Dieter Pfoser. ReHub: Extending hub labels for reverse k-nearest neighbor queries on large-scale networks. ACM Journal of Experimental Algorithmics, 21(1): 1.13:1–1.13:??, November 2016. CODEN ???? ISSN 1084-6654.

Efentakis:2017:REH

[EP16b] Alexandros Efentakis and Dieter Pfoser. ReHub: Extending hub labels for reverse k-nearest neighbor queries on large-scale networks. ACM Journal of Experimental Algorithmics, 21(1): 1.13:1–1.13:??, 2016. CODEN ???? ISSN 1084-6654.

Eppstein:2000:FHC

[Epp00] David Eppstein. Fast hierarchical clustering and other applications of dynamic closest pairs. ACM Journal of Experimental Algorithmics, 5:1:1–1:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Eiron:1999:MMC

[ERS99] N. Eiron, M. Rodeh, and I. Steinwarts. Matrix multiplication: a case study of enhanced data cache utilization. ACM Journal of Experimental Algorithmics, 4:3:1–3:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Englert:2009:EOS

[ERW09] Matthias Englert, Heiko Röglin, and Matthias Westermann. Evaluation of online strategies for reordering buffers. ACM Journal of Experimental Algorithmics, 14(1):3:1–3:??, December 2009. CODEN ???? ISSN 1084-6654.

Edelkamp:2002:IHQ

[ES02] Stefan Edelkamp and Patrick Stiegeler. Implementing HEAP-SORT with (n log n − 0.9n) and QUICKSORT with (n log n + 0.2n) comparisons. ACM Journal of Experimental Algorithmics, 7:5, ???? 2002. CODEN ???? ISSN 1084-6654.

Edelkamp:2019:BAB

[EW19] Stefan Edelkamp and Armin Weiß. BlockQuicksort: Avoiding branch mispredictions in Quicksort. ACM Journal of Experimental Algorithmics, 24(1):1.4:1–1.4:??, October 2019. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/ft_

gateway.cfm?id=3274660.

Farach-Colton:2021:DWS

[FCLMC21] Martín Farach-Colton, Katia Leal, Miguel A. Mosteiro, and Christopher Thraves Caro. Dynamic windows scheduling with reallocation. ACM Journal of Experimental Algorithmics, 26 (1):1–19, December 2021. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3462208.

Figueroa:2009:SSA

[FCNP09] Karina Figueroa, Edgar Chavez, Gonzalo Navarro, and Rodrigo Paredes. Speeding up spatial approximation search in metric spaces. ACM Journal of Experimental Algorithmics, 14(1):6:1– 6:??, December 2009. CODEN ???? ISSN 1084-6654.

Fox-Epstein:2016:SSC

[FEMPS16a] Eli Fox-Epstein, Shay Mozes, Phitchaya Mangpo Phothilimthana, and Christian Sommer. Short and simple cycle separators in planar graphs. ACM Journal of Experimental Algorithmics, 21(1): 2.2:1–2.2:??, November 2016. CODEN ???? ISSN 1084-6654.

Fox-Epstein:2017:SSC

[FEMPS16b] Eli Fox-Epstein, Shay Mozes, Phitchaya Mangpo Phothilimthana, and Christian Sommer. Short and simple cycle separators in planar graphs. ACM Journal of Experimental Algorithmics, 21(1): 2.2:1–2.2:??, 2016. CODEN ???? ISSN 1084-6654.

Festa:2013:FSI

[Fes13] Paola Festa. Foreword to the special issue SEA 2010. ACM Journal of Experimental Algorithmics, 18(1):1.1:1–1.1:??, December 2013. CODEN ???? ISSN 1084-6654.

Finocchi:2015:CCM

[FFF15] Irene Finocchi, Marco Finocchi, and Emanuele G. Fusco. Clique counting in MapReduce: Algorithms and experiments. ACM Journal of Experimental Algorithmics, 20(??):1.7:1–1.7:??, 2015. CODEN ???? ISSN 1084-6654.

Farshi:2009:ESG

[FG09] Mohammad Farshi and Joachim Gudmundsson. Experimental study of geometric t-spanners. ACM Journal of Experimental Algorithmics, 14(1):3:1–3:??, December 2009. CODEN ???? ISSN 1084-6654.

Fischl:2021:HBT

[FGLP21] Wolfgang Fischl, Georg Gottlob, Davide Mario Longo, and Reinhard Pichler. HyperBench: a benchmark and tool for hypergraphs and empirical findings. ACM Journal of Experimental Algorithmics, 26(1):1–40, December 2021. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/doi/10.1145/3440015.

Finocchi:2011:GEF

[FH11] Irene Finocchi and John Hershberger. Guest editors’ foreword. ACM Journal of Experimental Algorithmics, 16(1):31:1–31:??, 2011. CODEN ???? ISSN 1084-6654.

Flato:2000:DIP

[FHH^{+}00] Eyal Flato, Dan Halperin, Iddo Hanniel, Oren Nechushtan, and Eti Ezra. The design and implementation of planar maps in CGAL. ACM Journal of Experimental Algorithmics, 5:13:1– 13:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Fichte:2021:MCC

[FHH21] Johannes K. Fichte, Markus Hecher, and Florim Hamiti. The model counting competition 2020. ACM Journal of Experimental Algorithmics, 26(1):13:1–13:26, December 2021. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3459080.

Frigioni:1998:EAD

[FINP98] Daniele Frigioni, Mario Ioffreda, Umberto Nanni, and Giulio Pasqualone. Experimental analysis of dynamic algorithms for the single. ACM Journal of Experimental Algorithmics, 3:5:1– 5:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Fekete:2023:MPP

[FKKS23] Sándor P. Fekete, Phillip Keldenich, Dominik Krupke, and Stefan Schirra. Minimum partition into plane subgraphs: The CG:SHOP Challenge 2022. ACM Journal of Experimental Algorithmics, 28:1.9:1–1.9:??, 2023. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3604907.

Focke:2020:MST

[FMM20] Jacob Focke, Nicole Megow, and Julie Meißner. Minimum spanning tree under explorable uncertainty in theory and experiments. ACM Journal of Experimental Algorithmics, 25(1):1–20, April 2020. CODEN ???? ISSN 1084-6654. URL https://dl.

acm.org/doi/10.1145/3422371.

Fekete:2002:SHP

[FMRT02] Sándor P. Fekete, Henk Meijer, André Rohe, and Walter Tietze. Solving a ’hard’ problem to approximate an ’easy’ one: heuristics for maximum matchings and maximum traveling salesman problems. ACM Journal of Experimental Algorithmics, 7:11, ???? 2002. CODEN ???? ISSN 1084-6654.

Frigioni:2001:ESD

[FMZ01] Daniele Frigioni, Tobias Miller, and Christos Zaroliagis. An experimental study of dynamic algorithms for transitive closure. ACM Journal of Experimental Algorithmics, 6:9:1–9:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Fredriksson:2004:AOS

[FN04] Kimmo Fredriksson and Gonzalo Navarro. Average-optimal single and multiple approximate string matching. ACM Journal of Experimental Algorithmics, 9:1.4:1–1.4:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Flammini:2006:RAF

[FNP06] Michele Flammini, Alfredo Navarra, and Stephane Perennes. The “real” approximation factor of the MST heuristic for the minimum energy broadcasting. ACM Journal of Experimental Algorithmics, 11:2.10:1–2.10:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Ferraro-Petrillo:2013:DSR

[FPGI13] Umberto Ferraro-Petrillo, Fabrizio Grandoni, and Giuseppe F. Italiano. Data structures resilient to memory faults: an experimental study of dictionaries. ACM Journal of Experimental Algorithmics, 18(1):1.6:1–1.6:??, December 2013. CODEN ???? ISSN 1084-6654.

Festa:2006:GPR

[FPPR06] Paola Festa, Panos M. Pardalos, Leonidas S. Pitsoulis, and Mauricio G. C. Resende. GRASP with path relinking for the weighted MAXSAT problem. ACM Journal of Experimental Algorithmics, 11:2.4:1–2.4:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Frias:2009:LRC

[FPR09] Leonor Frias, Jordi Petit, and Salvador Roura. Lists revisited: Cache-conscious STL lists. ACM Journal of Experimental Algorithmics, 14(1):5:1–5:??, December 2009. CODEN ???? ISSN 1084-6654.

Fink:2023:ECP

[FPR23] Simon D. Fink, Matthias Pfretzschner, and Ignaz Rutter. Experimental comparison of PC-Trees and PQ-Trees. ACM Journal of Experimental Algorithmics, 28:1.10:1–1.10:??, 2023. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/doi/

10.1145/3611653.

Frieder:2015:ESA

[FR15] Asaf Frieder and Liam Roditty. An experimental study on approximating k shortest simple paths. ACM Journal of Experimental Algorithmics, 19(??):1.5:1–1.5:??, February 2015. CO-DEN ???? ISSN 1084-6654.

Faraj:2022:BSG

[FS22] Marcelo Fonseca Faraj and Christian Schulz. Buffered streaming graph partitioning. ACM Journal of Experimental Algorithmics, 27(1):1.10:1–1.10:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3546911.

Fahle:2006:FBB

[FT06] Torsten Fahle and Karsten Tiemann. A faster branch-and-bound algorithm for the test-cover problem based on setcovering techniques. ACM Journal of Experimental Algorithmics, 11:2.2:1–2.2:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Gonzalez:2020:WEA

[GC20] Miguel Coviello Gonzalez and Marek Chrobak. A waste-efficient algorithm for single-droplet sample preparation on microfluidic chips. ACM Journal of Experimental Algorithmics, 25(1):1–22, April 2020. CODEN ???? ISSN 1084-6654. URL https://dl.

acm.org/doi/10.1145/3408297.

Galvao:2015:ATG

[GD15] Gustavo Rodrigues Galvão and Zanoni Dias. An audit tool for genome rearrangement algorithms. ACM Journal of Experimental Algorithmics, 19(??):1.7:1–1.7:??, February 2015. CODEN ???? ISSN 1084-6654.

Gramm:2008:DRE

[GGHN08] Jens Gramm, Jiong Guo, Falk Hüffner, and Rolf Niedermeier. Data reduction and exact algorithms for clique cover. ACM Journal of Experimental Algorithmics, 13:2:1–2:??, September 2008. CODEN ???? ISSN 1084-6654.

Goodrich:2021:UEE

[GHS21] Timothy D. Goodrich, Eric Horton, and Blair D. Sullivan. An updated experimental evaluation of graph bipartization methods. ACM Journal of Experimental Algorithmics, 26(1):12:1– 12:24, December 2021. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3467968.

Gittleman:1996:PSS

[Git96] Arthur Gittleman. Predicting string search speed. ACM Journal of Experimental Algorithmics, 1:2:1–2:??, ???? 1996. CODEN ???? ISSN 1084-6654.

Grotschel:2022:CEM

[GJR22] Martin Grötschel, Michael Jünger, and Gerhard Reinelt. Comments on “An Exact Method for the Minimum Feedback Arc Set Problem”. ACM Journal of Experimental Algorithmics, 27 (1):1.3:1–1.3:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3545001. See [BSNA21].

Gabow:2001:NFB

[GK01] Harold Gabow and Tadayoshi Kohno. A network-flow-based scheduler: Design, performance history, and experimental analysis. ACM Journal of Experimental Algorithmics, 6:3:1–3:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Gudmundsson:2016:ESS

[GK16] Joachim Gudmundsson and Jyrki Katajainen. Editorial, SEA 2014 special issue. ACM Journal of Experimental Algorithmics, 21(1):1.1:1–1.1:??, November 2016. CODEN ???? ISSN 1084-6654.

Gog:2017:PCI

[GKN17] Simon Gog, Roberto Konow, and Gonzalo Navarro. Practical compact indexes for top-k document retrieval. ACM Journal of Experimental Algorithmics, 22(??):1.2:1–1.2:??, 2017. CODEN ???? ISSN 1084-6654.

Gorke:2015:EDC

[GKW15] Robert Görke, Andrea Kappes, and Dorothea Wagner. Experiments on density-constrained graph clustering. ACM Journal of Experimental Algorithmics, 19(??):1.6:1–1.6:??, February 2015. CODEN ???? ISSN 1084-6654.

Goshi:2006:ADM

[GL06] Justin Goshi and Richard E. Ladner. Algorithms for dynamic multicast key distribution. ACM Journal of Experimental Algorithmics, 11:1.4:1–1.4:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Graf:2020:XFF

[GL20] Thomas Mueller Graf and Daniel Lemire. Xor filters: Faster and smaller than Bloom and cuckoo filters. ACM Journal of Experimental Algorithmics, 25(1):1–16, April 2020. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/abs/
10.1145/3376122.

Graf:2022:BFF

[GL22] Thomas Mueller Graf and Daniel Lemire. Binary fuse filters: Fast and smaller than xor filters. ACM Journal of Experimental Algorithmics, 27(1):1.5:1–1.5:??, 2022. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/doi/10.1145/3510449.

Gottlob:2022:IUG

[GLLO22] Georg Gottlob, Matthias Lanzinger, Davide Mario Longo, and Cem Okulmus. Incremental updates of generalized hypertree decompositions. ACM Journal of Experimental Algorithmics, 27(1):1.16:1–1.16:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3578266.

Gorke:2013:DGC

[GMS^{+}13] Robert Görke, Pascal Maillard, Andrea Schumm, Christian Staudt, and Dorothea Wagner. Dynamic graph clustering combining modularity and smoothness. ACM Journal of Experimental Algorithmics, 18(1):1.5:1–1.5:??, December 2013. CO-DEN ???? ISSN 1084-6654.

Glantz:2016:TBC

[GMS16] Roland Glantz, Henning Meyerhenke, and Christian Schulz. Tree-based coarsening and partitioning of complex networks. ACM Journal of Experimental Algorithmics, 21(1):1.6:1–1.6:??, November 2016. CODEN ???? ISSN 1084-6654.

Gonzalez:2015:LCS

[GNF15] Rodrigo González, Gonzalo Navarro, and Héctor Ferrada. Locally compressed suffix arrays. ACM Journal of Experimental Algorithmics, 19(??):1.1:1–1.1:??, February 2015. CODEN ???? ISSN 1084-6654.

Gemsa:2016:ELS

[GNR16] Andreas Gemsa, Martin Nöllenburg, and Ignaz Rutter. Evaluation of labeling strategies for rotating maps. ACM Journal of Experimental Algorithmics, 21(1):1.4:1–1.4:??, November 2016. CODEN ???? ISSN 1084-6654.

Gog:2013:CST

[GO13] Simon Gog and Enno Ohlebusch. Compressed suffix trees: Efficient computation and storage of LCP-values. ACM Journal of Experimental Algorithmics, 18(1):2.1:1–2.1:??, December 2013. CODEN ???? ISSN 1084-6654.

Grossi:2015:FCT

[GO15] Roberto Grossi and Giuseppe Ottaviano. Fast compressed tries through path decompositions. ACM Journal of Experimental Algorithmics, 19(??):1.8:1–1.8:??, February 2015. CODEN ???? ISSN 1084-6654.

Geisberger:2012:RPF

[GRST12] Robert Geisberger, Michael N. Rice, Peter Sanders, and Vassilis J. Tsotras. Route planning with flexible edge restrictions. ACM Journal of Experimental Algorithmics, 17(1):1.2:1–1.2:??, March 2012. CODEN ???? ISSN 1084-6654.

Gottlob:2008:BBA

[GS08] Georg Gottlob and Marko Samer. A backtracking-based algorithm for hypertree decomposition. ACM Journal of Experimental Algorithmics, 13:1:1–1:??, September 2008. CODEN ???? ISSN 1084-6654.

Goh:2004:TAP

[GT04] Rick Siow Mong Goh and Ian Li-Jin Thng. Twol-amalgamated priority queues. ACM Journal of Experimental Algorithmics, 9: 1.6:1–1.6:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Hedtke:2015:UST

[Hed15] Ivo Hedtke. Upgrading subgroup triple-product-property triples. ACM Journal of Experimental Algorithmics, 20(??): 1.1:1–1.1:??, 2015. CODEN ???? ISSN 1084-6654.

Hyyro:2005:IBP

[HFN05] Heikki Hyyrö, Kimmo Fredriksson, and Gonzalo Navarro. Increased bit-parallelism for approximate and multiple string matching. ACM Journal of Experimental Algorithmics, 10: 2.6:1–2.6:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Hecht:2021:TLF

[HGH21] Michael Hecht, Krzysztof Gonciarz, and Szabolcs Horvát. Tight localizations of feedback sets. ACM Journal of Experimental Algorithmics, 26(1):1–19, December 2021. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/doi/10.1145/3447652.

Herrmann:2002:FCN

[HH02] Francine Herrmann and Alain Hertz. Finding the chromatic number by means of critical graphs. ACM Journal of Experimental Algorithmics, 7:10, ???? 2002. CODEN ???? ISSN 1084-6654.

Haran:2008:ESP

[HH08] Idit Haran and Dan Halperin. An experimental study of point location in planar arrangements in CGAL. ACM Journal of Experimental Algorithmics, 13:3:1–3:??, September 2008. CODEN ???? ISSN 1084-6654.

Huber:2011:MGS

[HH11] Stefan Huber and Martin Held. Motorcycle graphs: Stochastic properties motivate an efficient yet simple implementation. ACM Journal of Experimental Algorithmics, 16(1):13:1–13:??, 2011. CODEN ???? ISSN 1084-6654.

Hanauer:2022:RAF

[HHS22] Kathrin Hanauer, Monika Henzinger, and Christian Schulz. Recent advances in fully dynamic graph algorithms — a quick reference guide. ACM Journal of Experimental Algorithmics, 27 (1):1.11:1–1.11:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3555806.

Helman:1998:NDP

[HJB98] David R. Helman, Joseph JáJá, and David A. Bader. A new deterministic parallel sorting algorithm with an experimental evaluation. ACM Journal of Experimental Algorithmics, 3:4:1– 4:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Heinrich-Litan:2006:RCR

[HLL06] Laura Heinrich-Litan and Marco E. Lübbecke. Rectangle covers revisited computationally. ACM Journal of Experimental Algorithmics, 11:2.6:1–2.6:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Hamann:2018:EGM

[HMP^{+}18] Michael Hamann, Ulrich Meyer, Manuel Penschuck, Hung Tran, and Dorothea Wagner. I/O-efficient generation of massive graphs following the LFR benchmark. ACM Journal of Experimental Algorithmics, 23(??):2.5:1–2.5:??, 2018. CODEN ???? ISSN 1084-6654.

Huson:1999:HTR

[HNR^{+}99] D. Huson, S. Nettles, K. Rice, T. Warnow, and S. Yooseph. Hybrid tree reconstruction methods. ACM Journal of Experimental Algorithmics, 4:5:1–5:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Henzinger:2020:IBL

[HNS20] Alexandra Henzinger, Alexander Noe, and Christian Schulz. ILP-based local search for graph partitioning. ACM Journal of Experimental Algorithmics, 25(1):1–26, April 2020. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3398634.

Henzinger:2018:PMC

[HNSS18] Monika Henzinger, Alexander Noe, Christian Schulz, and Darren Strash. Practical minimum cut algorithms. ACM Journal of Experimental Algorithmics, 23(??):1.8:1–1.8:??, 2018. CODEN ???? ISSN 1084-6654.

Hofri:2013:OSS

[Hof13] Micha Hofri. Optimal selection and sorting via dynamic programming. ACM Journal of Experimental Algorithmics, 18(1): 2.3:1–2.3:??, December 2013. CODEN ???? ISSN 1084-6654.

Hemaspaandra:1998:PBA

[HRSZ98] Lane A. Hemaspaandra, Kulathur S. Rajasethupathy, Prasanna Sethupathy, and Marius Zimand. Power balance and apportionment algorithms for the United States Congress. ACM Journal of Experimental Algorithmics, 3:1:1–1:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Hamann:2018:GBP

[HS18] Michael Hamann and Ben Strasser. Graph bisection with Pareto optimization. ACM Journal of Experimental Algorithmics, 23 (??):1.2:1–1.2:??, 2018. CODEN ???? ISSN 1084-6654.

Hershberger:2008:SSD

[HSS08] John Hershberger, Nisheeth Shrivastava, and Subhash Suri. Summarizing spatial data streams using ClusterHulls. ACM Journal of Experimental Algorithmics, 13:4:1–4:??, September 2008. CODEN ???? ISSN 1084-6654.

Hespe:2019:SKM

[HSS19a] Demian Hespe, Christian Schulz, and Darren Strash. Scalable kernelization for maximum independent sets. ACM Journal of Experimental Algorithmics, 24(1):1.16:1–1.16:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://dl.acm.org/

ft_gateway.cfm?id=3355502.

Heuer:2019:NFB

[HSS19b] Tobias Heuer, Peter Sanders, and Sebastian Schlag. Network flow-based refinement for multilevel hypergraph partitioning. ACM Journal of Experimental Algorithmics, 24(1):2.3:1–2.3:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/ft_gateway.cfm?id=3329872.

Holzer:2008:EMO

[HSW08] Martin Holzer, Frank Schulz, and Dorothea Wagner. Engineering multilevel overlay graphs for shortest-path queries. ACM Journal of Experimental Algorithmics, 13:5:1–5:??, September 2008. CODEN ???? ISSN 1084-6654.

Holzer:2009:EPS

[HSW^{+}09] Martin Holzer, Frank Schulz, Dorothea Wagner, Grigorios Prasinos, and Christos Zaroliagis. Engineering planar separator algorithms. ACM Journal of Experimental Algorithmics, 14(1): 5:1–5:??, December 2009. CODEN ???? ISSN 1084-6654.

Holzer:2005:CST

[HSWW05] Martin Holzer, Frank Schulz, Dorothea Wagner, and Thomas Willhalm. Combining speed-up techniques for shortest-path computations. ACM Journal of Experimental Algorithmics, 10: 2.5:1–2.5:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Hirvola:2017:BPA

[HT17] Tommi Hirvola and Jorma Tarhio. Bit-parallel approximate matching of circular strings with k mismatches. ACM Journal of Experimental Algorithmics, 22(??):1.5:1–1.5:??, 2017. CODEN ???? ISSN 1084-6654.

Hazel:2008:TCL

[HTVW08] Thomas Hazel, Laura Toma, Jan Vahrenhold, and Rajiv Wickremesinghe. Terracost: Computing least-cost-path surfaces for massive grid terrains. ACM Journal of Experimental Algorithmics, 12:1.9:1–1.9:??, June 2008. CODEN ???? ISSN 1084-6654.

Haverkort:2011:FDH

[HW11] Herman Haverkort and Freek V. Walderveen. Four-dimensional Hilbert curves for R-trees. ACM Journal of Experimental Algorithmics, 16(1):34:1–34:??, 2011. CODEN ???? ISSN 1084-6654.

Ioannidis:2005:ADS

[IGA05] Ioannis Ioannidis, Ananth Grama, and Mikhail Atallah. Adaptive data structures for IP lookups. ACM Journal of Experimental Algorithmics, 10:1.1:1–1.1:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Ilinkin:2006:HEC

[IJS^{+}06] Ivayio Ilinkin, Ravi Janardan, Michiel Smid, Eric Johnson, Paul Castillo, and Jörg Schwerdt. Heuristics for estimating contact area of supports in layered manufacturing. ACM Journal of Experimental Algorithmics, 11:1.6:1–1.6:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Iwama:2002:PLS

[IKM^{+}02] Kazuo Iwama, Daisuke Kawai, Shuichi Miyazaki, Yasuo Okabe, and Jun Umemoto. Parallelizing local search for CNF satisfiability using vectorization and PVM. ACM Journal of Experimental Algorithmics, 7:2, ???? 2002. CODEN ???? ISSN 1084-6654.

Iyer:2001:ESP

[IKRT01] Raj Iyer, David Karger, Hariharan Rahul, and Mikkel Thorup. An experimental study of polylogarithmic, fully dynamic, connectivity algorithms. ACM Journal of Experimental Algorithmics, 6:4:1–4:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Irving:2009:FLS

[IM09] Robert W. Irving and David F. Manlove. Finding large stable matchings. ACM Journal of Experimental Algorithmics, 14:2:1– 2:??, May 2009. CODEN ???? ISSN 1084-6654.

Jacobs:2010:ESR

[Jac10] Tobias Jacobs. An experimental study of recent hotlink assignment algorithms. ACM Journal of Experimental Algorithmics, 15(1):1.1:1–1.1:??, March 2010. CODEN ???? ISSN 1084-6654.

Jensen:2020:SIS

[JB20] Alathea Jensen and Isabel Beichl. A sequential importance sampling algorithm for counting linear extensions. ACM Journal of Experimental Algorithmics, 25(1):1–14, April 2020. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3385650.

Junger:2021:QAV

[JLM^{+}21] Michael Jünger, Elisabeth Lobe, Petra Mutzel, Gerhard Reinelt, Franz Rendl, Giovanni Rinaldi, and Tobias Stollenwerk. Quantum annealing versus digital computing: an experimental comparison. ACM Journal of Experimental Algorithmics, 26(1):1– 30, December 2021. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3459606.

Jurkiewicz:2015:MVA

[JM15] Tomasz Jurkiewicz and Kurt Mehlhorn. On a model of virtual address translation. ACM Journal of Experimental Algorithmics, 19(??):1.9:1–1.9:??, February 2015. CODEN ???? ISSN 1084-6654.

Jacob:1999:CSR

[JMN99] R. Jacob, M. Marathe, and K. Nagel. A computational study of routing algorithms for realistic transportation networks. ACM Journal of Experimental Algorithmics, 4:6:1–6:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Julstrom:2009:GHB

[Jul09] Bryant A. Julstrom. Greedy heuristics for the bounded diameter minimum spanning tree problem. ACM Journal of Experimental Algorithmics, 14:1:1–1:??, May 2009. CODEN ???? ISSN 1084-6654.

Korhonen:2019:SGP

[KBJ19] Tuukka Korhonen, Jeremias Berg, and Matti Järvisalo. Solving graph problems via potential maximal cliques: an experimental evaluation of the Bouchitté–Todinca algorithm. ACM Journal of Experimental Algorithmics, 24(1):1.9:1–1.9:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://dl.acm.org/

ft_gateway.cfm?id=3301297.

Kot:2011:ECP

[KCC11] Andriy Kot, Andrey N. Chernikov, and Nikos P. Chrisochoides. Effective out-of-core parallel Delaunay mesh refinement using off-the-shelf software. ACM Journal of Experimental Algorithmics, 16(1):15:1–15:??, 2011. CODEN ???? ISSN 1084-6654.

Kim:1999:NSP

[Kim99] Sun Kim. A new string-pattern matching algorithm using partitioning and hashing efficiently. ACM Journal of Experimental Algorithmics, 4:2:1–2:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Karkkainen:2016:LAC

[KK16] Juha Kärkkäinen and Dominik Kempa. LCP array construction in external memory. ACM Journal of Experimental Algorithmics, 21(1):1.7:1–1.7:??, November 2016. CODEN ???? ISSN 1084-6654.

Karkkainen:2019:BEM

[KK19] Juha Kärkkäinen and Dominik Kempa. Better external memory LCP array construction. ACM Journal of Experimental Algorithmics, 24(1):1.3:1–1.3:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3297723.

Koana:2021:DRM

[KKN^{+}21] Tomohiro Koana, Viatcheslav Korenwein, André Nichterlein, Rolf Niedermeier, and Philipp Zschoche. Data reduction for maximum matching on real-world graphs: Theory and experiments. ACM Journal of Experimental Algorithmics, 26(1):1–30, December 2021. CODEN ???? ISSN 1084-6654. URL https:/

/dl.acm.org/doi/10.1145/3439801.

Karkkainen:2016:LLZ

[KKP16a] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, November 2016. CO-DEN ???? ISSN 1084-6654.

Karkkainen:2017:LLZ

[KKP16b] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, 2016. CODEN ???? ISSN 1084-6654.

Kanda:2020:DPD

[KKT^{+}20] Shunsuke Kanda, Dominik Köppl, Yasuo Tabei, Kazuhiro Morita, and Masao Fuketa. Dynamic path-decomposed tries. ACM Journal of Experimental Algorithmics, 25(1):1–28, April 2020. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/doi/10.1145/3418033.

Klasing:2015:E

[Kla15] Ralf Klasing. Editorial. ACM Journal of Experimental Algorithmics, 19(??):2.1:1–2.1:??, February 2015. CODEN ???? ISSN 1084-6654.

Kirchler:2015:ECS

[KLC15] Dominik Kirchler, Leo Liberti, and Roberto Wolfler Calvo. Efficient computation of shortest paths in time-dependent multimodal networks. ACM Journal of Experimental Algorithmics, 19(??):2.5:1–2.5:??, February 2015. CODEN ???? ISSN 1084-

Kouri:2013:FRM

[KM13a] Tina M. Kouri and Dinesh P. Mehta. Faster reaction mapping through improved naming techniques. ACM Journal of Experimental Algorithmics, 18(1):2.5:1–2.5:??, December 2013. CODEN ???? ISSN 1084-6654.

Kouzinopoulos:2013:EOT

[KM13b] Charalampos S. Kouzinopoulos and Konstantinos G. Margaritis. Exact online two-dimensional pattern matching using multiple pattern matching algorithms. ACM Journal of Experimental Algorithmics, 18(1):2.4:1–2.4:??, December 2013. CODEN ???? ISSN 1084-6654.

Khuong:2017:ALC

[KM17] Paul-Virak Khuong and Pat Morin. Array layouts for comparison-based searching. ACM Journal of Experimental Algorithmics, 22(??):1.3:1–1.3:??, 2017. CODEN ???? ISSN 1084-6654.

Kerber:2017:GHC

[KMN17] Michael Kerber, Dmitriy Morozov, and Arnur Nigmetov. Geometry helps to compare persistence diagrams. ACM Journal of Experimental Algorithmics, 22(??):1.4:1–1.4:??, 2017. CODEN ???? ISSN 1084-6654.

Kumar:2003:AME

[KMY03] Piyush Kumar, Joseph S. B. Mitchell, and E. Alper Yildirim. Approximate minimum enclosing balls in high dimensions using core-sets. ACM Journal of Experimental Algorithmics, 8:1.1:1– 1.1:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Knuth:1996:II

[Knu96] Donald E. Knuth. Irredundant intervals. ACM Journal of Experimental Algorithmics, 1:1:1–1:??, ???? 1996. CODEN ???? ISSN 1084-6654.

Kirchbach:2020:BPM

[KST20] Konrad Von Kirchbach, Christian Schulz, and Jesper Larsson Träff. Better process mapping and sparse quadratic assignment. ACM Journal of Experimental Algorithmics, 25(1):1–19, April 2020. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/doi/10.1145/3409667.

Krommidas:2008:ESA

[KZ08] Ioannis Krommidas and Christos Zaroliagis. An experimental study of algorithms for fully dynamic transitive closure. ACM Journal of Experimental Algorithmics, 12:16:1–16:??, June 2008. CODEN ???? ISSN 1084-6654.

Levine:2000:FRC

[Lev00] Matthew S. Levine. Finding the right cutting planes for the TSP. ACM Journal of Experimental Algorithmics, 5:6:1–6:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Levin:2019:ACR

[LF19] Harry A. Levin and Sorelle A. Friedler. Automated Congressional redistricting. ACM Journal of Experimental Algorithmics, 24(1):1.10:1–1.10:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3316513.

Leaver-Fay:2008:FPH

[LFLSW08] Andrew Leaver-Fay, Yuanxin Liu, Jack Snoeyink, and Xueyi Wang. Faster placement of hydrogens in protein structures by dynamic programming. ACM Journal of Experimental Algorithmics, 12:2.5:1–2.5:??, June 2008. CODEN ???? ISSN 1084-6654.

Lassous:2002:PLR

[LG02] Isabelle Guérin Lassous and Jens Gustedt. Portable list ranking: an experimental study. ACM Journal of Experimental Algorithmics, 7:7, ???? 2002. CODEN ???? ISSN 1084-6654.

Li:2008:EAC

[Li08] Keqin Li. Experimental average-case performance evaluation of online algorithms for routing and wavelength assignment and throughput maximization in WDM optical networks. ACM Journal of Experimental Algorithmics, 12:1.7:1–1.7:??, June 2008. CODEN ???? ISSN 1084-6654.

Liberatore:2001:CSB

[Lib01] Vincenzo Liberatore. Caching and scheduling for broadcast disk systems. ACM Journal of Experimental Algorithmics, 6:5:1– 5:??, ???? 2001. CODEN ???? ISSN 1084-6654.

LaMarca:1996:ICP

[LL96] Anthony LaMarca and Richard Ladner. The influence of caches on the performance of heaps. ACM Journal of Experimental Algorithmics, 1:4:1–4:??, ???? 1996. CODEN ???? ISSN 1084-6654.

Liu:2022:DOL

[LLH^{+}22] Xu Liu, Andrew Lumsdaine, Mahantesh Halappanavar, Kevin Barker, and Assefaw Gebremedhin. Direction-optimizing label propagation framework for structure detection in graphs: Design, implementation, and experimental analysis. ACM Journal of Experimental Algorithmics, 27(1):1.12:1–1.12:??, 2022. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/doi/

10.1145/3564593.

Lesh:2005:NHI

[LMMM05] N. Lesh, J. Marks, A. McMahon, and M. Mitzenmacher. New heuristic and interactive approaches to 2D rectangular strip packing. ACM Journal of Experimental Algorithmics, 10:1.2:1– 1.2:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Liberti:2023:RPL

[LMP23] Leo Liberti, Benedetto Manca, and Pierre-Louis Poirion. Random projections for linear programming: an improved retrieval phase. ACM Journal of Experimental Algorithmics, 28:2.2:1– 2.2:??, 2023. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/doi/10.1145/3617506.

Langguth:2010:HIB

[LMS10] Johannes Langguth, Fredrik Manne, and Peter Sanders. Heuristic initialization for bipartite matching problems. ACM Journal of Experimental Algorithmics, 15(1):1.3:1–1.3:??, March 2010. CODEN ???? ISSN 1084-6654.

Lanthier:2008:CAC

[LNW08] Mark A. Lanthier, Doron Nussbaum, and Tsuo-Jung Wang. Computing an approximation of the 1-center problem on weighted terrain surfaces. ACM Journal of Experimental Algorithmics, 13:3:1–3:??, September 2008. CODEN ???? ISSN 1084-6654.

Lopez-Ortiz:2005:FSS

[LOMSS05] Alejandro López-Ortiz, Mehdi Mirzazadeh, Mohammad Ali Safari, and Hossein Sheikhattar. Fast string sorting using orderpreserving compression. ACM Journal of Experimental Algorithmics, 10:1.4:1–1.4:??, ???? 2005. CODEN ???? ISSN 1084-

Liptak:2021:PDC

[LPR21] Zsuzsanna Lipták, Simon J. Puglisi, and Massimiliano Rossi. Pattern discovery in colored strings. ACM Journal of Experimental Algorithmics, 26(1):1–26, December 2021. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3429280.

Leone:2006:FPN

[LRAM06] Pierre Leone, Jose Rolim, Paul Albuquerque, and Christian Mazza. A framework for probabilistic numerical evaluation of sensor networks: a case study of a localization protocol. ACM Journal of Experimental Algorithmics, 11:2.3:1–2.3:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Luxen:2015:CSA

[LS15] Dennis Luxen and Dennis Schieferdecker. Candidate sets for alternative routes in road networks. ACM Journal of Experimental Algorithmics, 19(??):2.7:1–2.7:??, February 2015. CODEN ???? ISSN 1084-6654.

Lorenz:2022:TUL

[LW22] Jan-Hendrik Lorenz and Florian Wörz. Toward an understanding of long-tailed runtimes of SLS algorithms. ACM Journal of Experimental Algorithmics, 27(1):1.14:1–1.14:??, 2022. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/doi/

10.1145/3569170.

Magun:1998:GMA

[Mag98] Jakob Magun. Greeding matching algorithms, an experimental study. ACM Journal of Experimental Algorithmics, 3:6:1–6:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Mann:2018:CCR

[Man18] Zoltán Ádám Mann. Complexity of coloring random graphs: an experimental study of the hardest region. ACM Journal of Experimental Algorithmics, 23(??):1.3:1–1.3:??, 2018. CODEN ???? ISSN 1084-6654.

McGeoch:2010:P

[McG10] Catherine C. McGeoch. Preface. ACM Journal of Experimental Algorithmics, 15(1):2.1:1–2.1:??, March 2010. CODEN ???? ISSN 1084-6654.

Mcgeoch:2016:MDJ

[Mcg16] Catherine Mcgeoch. In memoriam: David S. Johnson. ACM Journal of Experimental Algorithmics, 21(1):1.1:1–1.1:??, November 2016. CODEN ???? ISSN 1084-6654.

Muller-Hannemann:1999:IWM

[MHS99] M. Müller-Hannemann and A. Schwartz. Implementing weighted b-matching algorithms: towards a flexible software design. ACM Journal of Experimental Algorithmics, 4:7:1–7:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Muller-Hannemann:2000:IWM

[MHS00] Matthias Müller-Hannemann and Alexander Schwartz. Implementing weighted b-matching algorithms: insights from a computational study. ACM Journal of Experimental Algorithmics, 5:8:1–8:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Michail:2011:ECS

[Mic11] Dimitrios Michail. An experimental comparison of single-sided preference matching algorithms. ACM Journal of Experimental Algorithmics, 16(1):14:1–14:??, 2011. CODEN ???? ISSN 1084-6654.

Mehlhorn:2006:IMC

[MM06] Kurt Mehlhorn and Dimitrios Michail. Implementing minimum cycle basis algorithms. ACM Journal of Experimental Algorithmics, 11:2.5:1–2.5:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Marinov:2016:PAF

[MNG16] Martin Marinov, Nicholas Nash, and David Gregg. Practical algorithms for finding extremal sets. ACM Journal of Experimental Algorithmics, 21(1):1.9:1–1.9:??, November 2016. CODEN ???? ISSN 1084-6654.

Moruz:2015:EEP

[MNNW15] Gabriel Moruz, Andrei Negoescu, Christian Neumann, and Volker Weichert. Engineering efficient paging algorithms. ACM Journal of Experimental Algorithmics, 19(??):2.4:1– 2.4:??, February 2015. CODEN ???? ISSN 1084-6654.

Manlove:2015:PAK

[MO15] David F. Manlove and Gregg O’Malley. Paired and altruistic kidney donation in the UK: Algorithms and experimentation. ACM Journal of Experimental Algorithmics, 19(??):2.6:1– 2.6:??, February 2015. CODEN ???? ISSN 1084-6654.

Maniscalco:2008:EVA

[MP08] Michael A. Maniscalco and Simon J. Puglisi. An efficient, versatile approach to suffix sorting. ACM Journal of Experimental Algorithmics, 12:1.2:1–1.2:??, June 2008. CODEN ???? ISSN 1084-6654.

Marathe:2004:ESS

[MPR04] Madhav V. Marathe, Alessandro Panconesi, and Larry D. Risinger, Jr. An experimental study of a simple, distributed edge-coloring algorithm. ACM Journal of Experimental Algorithmics, 9:1.3:1–1.3:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Mitzenmacher:2020:ACF

[MPR20] Michael Mitzenmacher, Salvatore Pontarelli, and Pedro Reviriego. Adaptive cuckoo filters. ACM Journal of Experimental Algorithmics, 25(1):1–20, April 2020. CODEN ???? ISSN 1084-

6654. URL https://dl.acm.org/doi/abs/10.1145/3339504.

Matias:2001:EFP

[MRS01] Yossi Matias, Nasir Rajpoot, and Cenk Sahinalp. The effect of flexible parsing for dynamic dictionary-based data compression. ACM Journal of Experimental Algorithmics, 6:10:1–10:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Mehlhorn:2002:IWM

[MS02] Kurt Mehlhorn and Guido Schäfer. Implementation of O(nm log n) weighted matchings in general graphs: the power of data structures. ACM Journal of Experimental Algorithmics, 7:4, ???? 2002. CODEN ???? ISSN 1084-6654.

Maue:2009:GDS

[MSM09] Jens Maue, Peter Sanders, and Domagoj Matijevic. Goaldirected shortest-path queries using precomputed cluster distances. ACM Journal of Experimental Algorithmics, 14(1):2:1– 2:??, December 2009. CODEN ???? ISSN 1084-6654.

Mohring:2006:PGS

[MSS^{+}06] Rolf H. Möhring, Heiko Schilling, Birk Schütz, Dorothea Wagner, and Thomas Willhalm. Partitioning graphs to speedup Dijkstra’s algorithm. ACM Journal of Experimental Algorithmics, 11:2.8:1–2.8:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Munro:2009:PSS

[MW09] J. Ian Munro and Dorothea Wagner. Preface: Section 2 — selected papers from ALENEX 2008. ACM Journal of Experimental Algorithmics, 14:1:1–1:??, May 2009. CODEN ???? ISSN 1084-6654.

Mirka:2023:EES

[MW23] Renee Mirka and David P. Williamson. An experimental evaluation of semidefinite programming and spectral algorithms for max cut. ACM Journal of Experimental Algorithmics, 28:2.1:1– 2.1:??, 2023. CODEN ???? ISSN 1084-6654. URL https://

dl.acm.org/doi/10.1145/3609426.

Niewiadomski:2004:PSD

[NAH04] Robert Niewiadomski, José Nelson Amaral, and Robert C. Holte. A performance study of data layout techniques for improving data locality in refinement-based pathfinding. ACM Journal of Experimental Algorithmics, 9:1.2:1–1.2:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Neri:2002:RCL

[Ner02] Filippo Neri. Relational concept learning by cooperative evolution. ACM Journal of Experimental Algorithmics, 7:12, ???? 2002. CODEN ???? ISSN 1084-6654.

Nash:2010:CID

[NG10] Nicholas Nash and David Gregg. Comparing integer data structures for 32- and 64-bit keys. ACM Journal of Experimental Algorithmics, 15(1):2.4:1–2.4:??, March 2010. CODEN ???? ISSN 1084-6654.

Netto:2022:SSA

[NG22] Marcelo Vaz Netto and Sahudy Montenegro González. SSLC: a search algorithm based on linear collisions and Poisson probability distribution. ACM Journal of Experimental Algorithmics, 27(1):1.4:1–1.4:??, 2022. CODEN ???? ISSN 1084-6654. URL

https://dl.acm.org/doi/10.1145/3497876.

Nikoletseas:2006:JSS

[Nik06] Sotiris Nikoletseas. JEA Special Section. ACM Journal of Experimental Algorithmics, 11:2.1:1–2.1:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Nunes:2022:GCI

[NLG^{+}22] Daniel S. N. Nunes, Felipe A. Louza, Simon Gog, Mauricio Ayala-Rincón, and Gonzalo Navarro. Grammar compression by induced suffix sorting. ACM Journal of Experimental Algorithmics, 27(1):1.1:1–1.1:??, 2022. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3549992.

Navarro:2016:FCS

[NP16] Gonzalo Navarro and Alberto Ordóñez Pereira. Faster compressed suffix trees for repetitive collections. ACM Journal of Experimental Algorithmics, 21(1):1.8:1–1.8:??, November 2016. CODEN ???? ISSN 1084-6654.

Nadara:2019:EEA

[NPR^{+}19] Wojciech Nadara, Marcin Pilipczuk, Roman Rabinovich, Felix Reidl, and Sebastian Siebertz. Empirical evaluation of approximation algorithms for generalized graph coloring and uniform quasi-wideness. ACM Journal of Experimental Algorithmics, 24 (2):2.6:1–2.6:??, December 2019. CODEN ???? ISSN 1084-6654.

Negrucseri:2011:SMF

[NPS^{+}11] Cosmin Silvestru Negrucseri, Mircea Bogdan Pacsosi, Barbara Stanley, Clifford Stein, and Cristian George Strat. Solving maximum flow problems on real-world bipartite graphs. ACM Journal of Experimental Algorithmics, 16(1):35:1–35:??, 2011. CO-DEN ???? ISSN 1084-6654.

Navarro:2015:GDR

[NPV15] Gonzalo Navarro, Simon J. Puglisi, and Daniel Valenzuela. General document retrieval in compact space. ACM Journal of Experimental Algorithmics, 19(??):2.3:1–2.3:??, February 2015. CODEN ???? ISSN 1084-6654.

Navarro:2000:FFS

[NR00] Gonzalo Navarro and Mathieu Raffinot. Fast and flexible string matching by combining bit-parallelism and suffix automata. ACM Journal of Experimental Algorithmics, 5:4:1–4:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Navarro:2008:DSA

[NR08] Gonzalo Navarro and Nora Reyes. Dynamic spatial approximation trees. ACM Journal of Experimental Algorithmics, 12: 1.5:1–1.5:??, June 2008. CODEN ???? ISSN 1084-6654.

Nikolov:2005:SEH

[NTB05] Nikola S. Nikolov, Alexandre Tarassov, and Jürgen Branke. In search for efficient heuristics for minimum-width graph layering with consideration of dummy nodes. ACM Journal of Experimental Algorithmics, 10:2.7:1–2.7:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Nagarajan:2013:EEI

[NW13] Chandrashekhar Nagarajan and David P. Williamson. An experimental evaluation of incremental and hierarchical k-median algorithms. ACM Journal of Experimental Algorithmics, 18(??): 3.2:1–3.2:??, December 2013. CODEN ???? ISSN 1084-6654.

Narasimhan:2001:GMS

[NZ01] Giri Narasimhan and Martin Zachariasen. Geometric minimum spanning trees via well-separated pair decompositions. ACM Journal of Experimental Algorithmics, 6:6:1–6:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Polischchuk:2018:EAS

[PC18] Valentin Polischchuk and Vijaya Ramachandran / Rezaul A. Chowdhury. Editorial: ALENEX 2017 special issue. ACM Journal of Experimental Algorithmics, 23(??):2.1:1–2.1:??, 2018. CODEN ???? ISSN 1084-6654.

Purchase:1997:ESB

[PCJ97] H. C. Purchase, R. F. Cohen, and M. I. James. An experimental study of the basis for graph drawing algorithms. ACM Journal of Experimental Algorithmics, 2:4:1–4:??, ???? 1997. CODEN ???? ISSN 1084-6654.

Petit:2003:EML

[Pet03] Jordi Petit. Experiments on the minimum linear arrangement problem. ACM Journal of Experimental Algorithmics, 8:2.3:1– 2.3:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Pellegrini:2008:EIT

[PF08] Marco Pellegrini and Giordano Fusco. Efficient IP table lookup via adaptive stratified trees with selective reconstructions. ACM Journal of Experimental Algorithmics, 12:1.4:1–1.4:??, June 2008. CODEN ???? ISSN 1084-6654.

Paudel:2018:CCN

[PGI18] Nilakantha Paudel, Loukas Georgiadis, and Giuseppe F. Italiano. Computing critical nodes in directed graphs. ACM Journal of Experimental Algorithmics, 23(??):2.2:1–2.2:??, 2018. CO-DEN ???? ISSN 1084-6654.

Pearce:2006:DTS

[PK06] David J. Pearce and Paul H. J. Kelly. A dynamic topological sort algorithm for directed acyclic graphs. ACM Journal of Experimental Algorithmics, 11:1.7:1–1.7:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Penner:2006:CFI

[PP06] Michael Penner and Viktor K. Prasanna. Cache-friendly implementations of transitive closure. ACM Journal of Experimental Algorithmics, 11:1.3:1–1.3:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Peethambaran:2016:ESR

[PPM16] Jiju Peethambaran, Amal Dev Parakkat, and Ramanathan Muthuganapathy. An empirical study on randomized optimal area polygonization of planar point sets. ACM Journal of Experimental Algorithmics, 21(1):1.10:1–1.10:??, November 2016. CODEN ???? ISSN 1084-6654.

Pemmaraju:2005:AIC

[PPR05] Sriram V. Pemmaraju, Sriram Penumatcha, and Rajiv Raman. Approximating interval coloring and max-coloring in chordal graphs. ACM Journal of Experimental Algorithmics, 10:2.8:1– 2.8:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Panagopoulou:2006:APN

[PS06] Panagiota N. Panagopoulou and Paul G. Spirakis. Algorithms for pure Nash equilibria in weighted congestion games. ACM Journal of Experimental Algorithmics, 11:2.7:1–2.7:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Putze:2009:CHS

[PSS09] Felix Putze, Peter Sanders, and Johannes Singler. Cache-, hash-, and space-efficient Bloom filters. ACM Journal of Experimental Algorithmics, 14(1):4:1–4:??, December 2009. CODEN ???? ISSN 1084-6654.

Pyrga:2008:EMT

[PSWZ08] Evangelia Pyrga, Frank Schulz, Dorothea Wagner, and Christos Zaroliagis. Efficient models for timetable information in public transportation systems. ACM Journal of Experimental Algorithmics, 12:2.4:1–2.4:??, June 2008. CODEN ???? ISSN 1084-6654.

Prosser:2011:LDS

[PU11] Patrick Prosser and Chris Unsworth. Limited discrepancy search revisited. ACM Journal of Experimental Algorithmics, 16(1):16:1–16:??, 2011. CODEN ???? ISSN 1084-6654.

Poloczek:2017:EEF

[PW17] Matthias Poloczek and David P. Williamson. An experimental evaluation of fast approximation algorithms for the maximum satisfiability problem. ACM Journal of Experimental Algorithmics, 22(??):1.6:1–1.6:??, 2017. CODEN ???? ISSN 1084-6654.

Qiu:2021:EER

[QSL^{+}21] Zirou Qiu, Ruslan Shaydulin, Xiaoyuan Liu, Yuri Alexeev, Christopher S. Henry, and Ilya Safro. ELRUNA: Elimination rule-based network alignment. ACM Journal of Experimental Algorithmics, 26(1):1–32, December 2021. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/doi/10.1145/3450703.

Radzik:1998:IDT

[Rad98] Tomasz Radzik. Implementation of dynamic trees with insubtree operations. ACM Journal of Experimental Algorithmics, 3:9:1–9:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Reams:2012:AFD

[Rea12] Charles Reams. Anatree: a fast data structure for anagrams. ACM Journal of Experimental Algorithmics, 17(1):1.1:1–1.1:??, March 2012. CODEN ???? ISSN 1084-6654.

Ribeiro:2005:P

[RM05] Celso C. Ribeiro and Simone L. Martins. Preface. ACM Journal of Experimental Algorithmics, 10:2.1:1–2.1:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Rosenbrock:2016:NAP

[RMH^{+}16] Conrad W. Rosenbrock, Wiley S. Morgan, Gus L. W. Hart, Stefano Curtarolo, and Rodney W. Forcade. Numerical algorithm for Pólya enumeration theorem. ACM Journal of Experimental Algorithmics, 21(1):1.11:1–1.11:??, November 2016. CODEN ???? ISSN 1084-6654.

Rotta:2011:MLS

[RN11] Randolf Rotta and Andreas Noack. Multilevel local search algorithms for modularity clustering. ACM Journal of Experimental Algorithmics, 16(1):23:1–23:??, 2011. CODEN ???? ISSN 1084-6654.

Rahman:2000:ACE

[RR00] Naila Rahman and Rajeev Raman. Analysing cache effects in distribution sorting. ACM Journal of Experimental Algorithmics, 5:14:1–14:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Rahman:2001:ARS

[RR01] Naila Rahman and Rajeev Raman. Adapting radix sort to the memory hierarchy. ACM Journal of Experimental Algorithmics, 6:7:1–7:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Radermacher:2019:GHR

[RRRW19] Marcel Radermacher, Klara Reichard, Ignaz Rutter, and Dorothea Wagner. Geometric heuristics for rectilinear crossing minimization. ACM Journal of Experimental Algorithmics, 24(1):1.12:1–1.12:??, October 2019. CODEN ???? ISSN

1084-6654. URL https://dl.acm.org/ft_gateway.cfm?id=
3325861.

Raman:2008:P

[RS08] Rajeev Raman and Matt Stallmann. Preface. ACM Journal of Experimental Algorithmics, 13:1:1–1:??, September 2008. CO-DEN ???? ISSN 1084-6654.

Robinson:2018:AHP

[RVHE18] Jeffrey A. Robinson, Susan V. Vrbsky, Xiaoyan Hong, and Brian P. Eddy. Analysis of a high-performance TSP solver on the GPU. ACM Journal of Experimental Algorithmics, 23(??): 1.1:1–1.1:??, 2018. CODEN ???? ISSN 1084-6654.

Serna:2009:PSS

[SÁ09] Maria Serna and Carme Álvarez. Preface to special section of selected papers from WEA 2006. ACM Journal of Experimental Algorithmics, 14(1):1:1–1:??, December 2009. CODEN ???? ISSN 1084-6654.

Sanders:2000:FPQ

[San00] Peter Sanders. Fast priority queues for cached memory. ACM Journal of Experimental Algorithmics, 5:7:1–7:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Stallmann:2001:HES

[SBG01] Matthias Stallmann, Franc Brglez, and Debabrata Ghosh. Heuristics, experimental subjects, and treatment evaluation in bigraph crossing minimization. ACM Journal of Experimental Algorithmics, 6:8:1–8:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Sieranoja:2018:CHD

[SF18] Sami Sieranoja and Pasi Fränti. Constructing a highdimensional k NN-graph using a Z-order curve. ACM Journal of Experimental Algorithmics, 23(??):1.9:1–1.9:??, 2018. CODEN ???? ISSN 1084-6654.

Stones:2020:CAG

[SFKM20] Rebecca J. Stones, Raúl M. Falcón, Daniel Kotlar, and Trent G. Marbach. Computing autotopism groups of partial Latin rectangles. ACM Journal of Experimental Algorithmics, 25(1):1–39, April 2020. CODEN ???? ISSN 1084-6654. URL https://dl.

acm.org/doi/10.1145/3412324.

Smith:1997:EHF

[SHA97] Bradley J. Smith, Gregory L. Heileman, and Chaouki Abdallah. The exponential hash function. ACM Journal of Experimental Algorithmics, 2:3:1–3:??, ???? 1997. CODEN ???? ISSN 1084-6654.

Schlag:2022:HQH

[SHG^{+}22] Sebastian Schlag, Tobias Heuer, Lars Gottesbüren, Yaroslav Akhremtsev, Christian Schulz, and Peter Sanders. High-quality hypergraph partitioning. ACM Journal of Experimental Algorithmics, 27(1):1.9:1–1.9:??, 2022. CODEN ???? ISSN 1084-

6654. URL https://dl.acm.org/doi/10.1145/3529090.

Shibuya:2000:CSP

[Shi00] Tetsuo Shibuya. Computing the n × m shortest path efficiently. ACM Journal of Experimental Algorithmics, 5:9:1–9:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Swenson:2008:ATE

[SMEDM08] Krister M. Swenson, Mark Marron, Joel V. Earnest-Deyoung, and Bernard M. E. Moret. Approximating the true evolutionary distance between two genomes. ACM Journal of Experimental Algorithmics, 12:3.5:1–3.5:??, June 2008. CODEN ???? ISSN 1084-6654.

Santos:2005:TSH

[SOS05] Haroldo G. Santos, Luiz S. Ochi, and Marcone J. F. Souza. A Tabu search heuristic with efficient diversification strategies for the class/teacher timetabling problem. ACM Journal of Experimental Algorithmics, 10:2.9:1–2.9:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Spence:2010:SGS

[Spe10] Ivor Spence. sgen1: a generator of small but difficult satisfiability benchmarks. ACM Journal of Experimental Algorithmics, 15(1):1.2:1–1.2:??, March 2010. CODEN ???? ISSN 1084-6654.

Spence:2015:WCC

[Spe15] Ivor Spence. Weakening cardinality constraints creates harder satisfiability benchmarks. ACM Journal of Experimental Algorithmics, 20(??):1.4:1–1.4:??, 2015. CODEN ???? ISSN 1084-6654.

Safro:2008:MAL

[SRB08] Ilya Safro, Dorit Ron, and Achi Brandt. Multilevel algorithms for linear ordering problems. ACM Journal of Experimental Algorithmics, 13:4:1–4:??, September 2008. CODEN ???? ISSN 1084-6654.

Schidler:2023:SBT

[SS23] André Schidler and Stefan Szeider. SAT-boosted tabu search for coloring massive graphs. ACM Journal of Experimental Algorithmics, 28:1.5:1–1.5:??, 2023. CODEN ???? ISSN 1084-6654.

URL https://dl.acm.org/doi/10.1145/3603112.

Schwerdt:1999:CWT

[SSMJ99] J. Schwerdt, M. Smid, J. Majhi, and R. Janardan. Computing the width of a three-dimensional point set: an experimental study. ACM Journal of Experimental Algorithmics, 4:8:1–8:??, ???? 1999. CODEN ???? ISSN 1084-6654.

Safro:2015:ACS

[SSS15] Ilya Safro, Peter Sanders, and Christian Schulz. Advanced coarsening schemes for graph partitioning. ACM Journal of Experimental Algorithmics, 19(??):2.2:1–2.2:??, February 2015. CODEN ???? ISSN 1084-6654.

Schlag:2021:FSV

[SSS21] Sebastian Schlag, Matthias Schmitt, and Christian Schulz. Faster support vector machines. ACM Journal of Experimental Algorithmics, 26(1):15:1–15:21, December 2021. CODEN

???? ISSN 1084-6654. URL https://dl.acm.org/doi/10.
1145/3484730.

Salmela:2006:MSM

[STK06] Leena Salmela, Jorma Tarhio, and Jari Kytöjoki. Multipattern string matching with q-grams. ACM Journal of Experimental Algorithmics, 11:1.1:1–1.1:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Stoichev:2019:NEH

[Sto19] Stoicho D. Stoichev. New exact and heuristic algorithms for graph automorphism group and graph isomorphism. ACM Journal of Experimental Algorithmics, 24(1):1.15:1–1.15:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://dl.acm.

org/ft_gateway.cfm?id=3333250.

Sinha:2010:EBT

[SW10] Ranjan Sinha and Anthony Wirth. Engineering burstsort: Toward fast in-place string sorting. ACM Journal of Experimental Algorithmics, 15(1):2.5:1–2.5:??, March 2010. CODEN ???? ISSN 1084-6654.

Schulz:2000:DAL

[SWW00] Frank Schulz, Dorothea Wagner, and Karsten Weihe. Dijkstra’s algorithm on-line: an empirical case study from public railroad transport. ACM Journal of Experimental Algorithmics, 5:12:1– 12:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Sinha:2004:CCS

[SZ04] Ranjan Sinha and Justin Zobel. Cache-conscious sorting of large sets of strings with dynamic tries. ACM Journal of Experimental Algorithmics, 9:1.5:1–1.5:??, ???? 2004. CODEN ???? ISSN 1084-6654.

Sinha:2005:URS

[SZ05] Ranjan Sinha and Justin Zobel. Using random sampling to build approximate tries for efficient string sorting. ACM Journal of Experimental Algorithmics, 10:2.10:1–2.10:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Sanders:2016:ISI

[SZ16a] Peter Sanders and Norbert Zeh. Introduction to special issue ALENEX 2013. ACM Journal of Experimental Algorithmics, 21(1):2.1:1–2.1:??, November 2016. CODEN ???? ISSN 1084-6654.

Sanders:2017:ISI

[SZ16b] Peter Sanders and Norbert Zeh. Introduction to special issue ALENEX 2013. ACM Journal of Experimental Algorithmics, 21 (1):2.1:1–2.1:??, 2016. CODEN ???? ISSN 1084-6654.

Sinha:2006:CES

[SZR06] Ranjan Sinha, Justin Zobel, and David Ring. Cache-efficient string sorting using copying. ACM Journal of Experimental Algorithmics, 11:1.2:1–1.2:??, ???? 2006. CODEN ???? ISSN 1084-6654.

Tazari:2011:DLH

[TMH11] Siamak Tazari and Matthias Müller-Hannemann. Dealing with large hidden constants: engineering a Planar Steiner Tree (PTAS). ACM Journal of Experimental Algorithmics, 16(1): 36:1–36:??, 2011. CODEN ???? ISSN 1084-6654.

Tsourakakis:2011:AAS

[TPT^{+}11] Charalampos E. Tsourakakis, Richard Peng, Maria A. Tsiarli, Gary L. Miller, and Russell Schwartz. Approximation algorithms for speeding up dynamic programming and denoising aCGH data. ACM Journal of Experimental Algorithmics, 16 (1):18:1–18:??, 2011. CODEN ???? ISSN 1084-6654.

Tabourier:2011:GCR

[TRC11] Lionel Tabourier, Camille Roth, and Jean-Philippe Cointet. Generating constrained random graphs using multiple edge switches. ACM Journal of Experimental Algorithmics, 16(1): 17:1–17:??, 2011. CODEN ???? ISSN 1084-6654.

Toda:2016:IEA

[TS16] Takahisa Toda and Takehide Soh. Implementing efficient all solutions SAT solvers. ACM Journal of Experimental Algorithmics, 21(1):1.12:1–1.12:??, November 2016. CODEN ???? ISSN 1084-6654.

Tsirogiannis:2018:CEV

[TSP18] Constantinos Tsirogiannis, Frank Staals, and Vincent Pellissier. Computing the expected value and variance of geometric measures. ACM Journal of Experimental Algorithmics, 23(??):2.4:1– 2.4:??, 2018. CODEN ???? ISSN 1084-6654.

Tarjan:2009:DTP

[TW09] Robert E. Tarjan and Renato F. Werneck. Dynamic trees in practice. ACM Journal of Experimental Algorithmics, 14(1): 5:1–5:??, December 2009. CODEN ???? ISSN 1084-6654.

Ullmann:2010:BVA

[Ull10] Julian R. Ullmann. Bit-vector algorithms for binary constraint satisfaction and subgraph isomorphism. ACM Journal of Experimental Algorithmics, 15(1):16:1–16:??, March 2010. CODEN ???? ISSN 1084-6654.

Ullmann:2015:DRL

[Ull15] Julian R. Ullmann. Degree reduction in labeled graph retrieval. ACM Journal of Experimental Algorithmics, 20(??): 1.3:1–1.3:??, 2015. CODEN ???? ISSN 1084-6654.

Vahrenhold:2011:P

[Vah11] Jan Vahrenhold. Preface. ACM Journal of Experimental Algorithmics, 16(1):21:1–21:??, 2011. CODEN ???? ISSN 1084-6654.

Vella:2018:DMF

[VBC18] Flavio Vella, Massimo Bernaschi, and Giancarlo Carbone. Dynamic merging of frontiers for accelerating the evaluation of betweenness centrality. ACM Journal of Experimental Algorithmics, 23(??):1.4:1–1.4:??, 2018. CODEN ???? ISSN 1084-6654.

vander:2022:SKR

[vdBG^{+}22] Alexander Grinten van der, Elisabetta Bergamini, Oded Green, David A. Bader, and Henning Meyerhenke. Scalable Katz ranking computation in large static and dynamic graphs. ACM Journal of Experimental Algorithmics, 27(1):1.7:1–1.7:??, 2022. CO-DEN ???? ISSN 1084-6654. URL https://dl.acm.org/doi/

10.1145/3524615.

Vahrenhold:2002:PPL

[VH02] Jan Vahrenhold and Klaus H. Hinrichs. Planar point location for large data sets: to seek or not to seek. ACM Journal of Experimental Algorithmics, 7:8, ???? 2002. CODEN ???? ISSN 1084-6654.

VonLooz:2018:UDR

[VM18] Moritz Von Looz and Henning Meyerhenke. Updating dynamic random hyperbolic graphs in sublinear time. ACM Journal of Experimental Algorithmics, 23(??):1.6:1–1.6:??, 2018. CODEN ???? ISSN 1084-6654.

Valimaki:2009:ECS

[VMGD09] N. Välimäki, V. Mäkinen, W. Gerlach, and K. Dixit. Engineering a compressed suffix tree implementation. ACM Journal of Experimental Algorithmics, 14(1):2:1–2:??, December 2009. CODEN ???? ISSN 1084-6654.

Venkataraman:2003:BAP

[VSM03] Gayathri Venkataraman, Sartaj Sahni, and Srabani Mukhopadhyaya. A blocked all-pairs shortest-paths algorithm. ACM Journal of Experimental Algorithmics, 8:2.2:1–2.2:??, ???? 2003. CODEN ???? ISSN 1084-6654.

Vishkin:2000:ELR

[VV00] Dascal Vishkin and Uzi Vishkin. Experiments with list ranking for explicit multi-threaded (XMT) instruction parallelism. ACM Journal of Experimental Algorithmics, 5:10:1–10:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Wickremesinghe:2002:ESU

[WACV02] Rajiv Wickremesinghe, Lars Arge, Jeffrey S. Chase, and Jeffrey Scott Vitter. Efficient sorting using registers and caches. ACM Journal of Experimental Algorithmics, 7:9, ???? 2002. CODEN ???? ISSN 1084-6654.

Wang:2011:CEM

[WEM11] Bei Wang, Herbert Edelsbrunner, and Dmitriy Morozov. Computing elevation maxima by searching the Gauss sphere. ACM Journal of Experimental Algorithmics, 16(1):22:1–22:??, 2011. CODEN ???? ISSN 1084-6654.

Werneck:2000:FMC

[WSdC00] Renato Werneck, João Setubal, and Arlindo da Conceicão. Finding minimum congestion spanning trees. ACM Journal of Experimental Algorithmics, 5:11:1–11:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Wagner:2005:GCE

[WWZ05] Dorothea Wagner, Thomas Willhalm, and Christos Zaroliagis. Geometric containers for efficient shortest-path computation. ACM Journal of Experimental Algorithmics, 10:1.3:1–1.3:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Xiao:2000:IMP

[XZK00] Li Xiao, Xiaodong Zhang, and Stefan A. Kubricht. Improving memory performance of sorting algorithms. ACM Journal of Experimental Algorithmics, 5:3:1–3:??, ???? 2000. CODEN ???? ISSN 1084-6654.

Yasar:2022:SRP

[YBA^{+}22] Abdurrahman Yaşar, Muhammed Fati˙ h Balin, Xiaojing An, Kaan Sancak, and Ümit V. Çatalyürek. On symmetric rectilinear partitioning. ACM Journal of Experimental Algorithmics, 27(1):1.2:1–1.2:??, 2022. CODEN ????

ISSN 1084-6654. URL https://dl.acm.org/doi/10.1145/
3523750.

Yan:1998:LBE

[YZ98] Yong Yan and Xiaodong Zhang. Lock bypassing: an efficient algorithm for concurrently accessing priority heaps. ACM Journal of Experimental Algorithmics, 3:3:1–3:??, ???? 1998. CODEN ???? ISSN 1084-6654.

Zaroliagis:2019:EES

[Zar19] Christos Zaroliagis. Editorial — ESA 2016 special issue. ACM Journal of Experimental Algorithmics, 24(1):1.1:1–1.1:??, October 2019. CODEN ???? ISSN 1084-6654. URL https://dl.

acm.org/ft_gateway.cfm?id=3298788.

Ziobro:2019:FHC

[ZP19] Michał Ziobro and Marcin Pilipczuk. Finding Hamiltonian cycle in graphs of bounded treewidth: Experimental evaluation. ACM Journal of Experimental Algorithmics, 24(2):2.7:1–2.7:18, December 2019. CODEN ???? ISSN 1084-6654.